### v1

# High-Temperature Fermionic Gibbs States are Mixtures of Gaussian States

Abstract
Efficient simulation of a quantum system generally relies on structural properties of the quantum state. Motivated by the recent results by Bakshi et al. on the sudden death of entanglement in high-temperature Gibbs states of quantum spin systems, we study the high-temperature Gibbs states of bounded-degree local fermionic Hamiltonians, which include the special case of geometrically local fermionic systems. We prove that at a sufficiently high temperature that is independent of the system size, the Gibbs state is a probabilistic mixture of fermionic Gaussian states. This forms the basis of an efficient classical algorithm to prepare the Gibbs state by sampling from a distribution of fermionic Gaussian states.
1 Introduction
Studying quantum many-body systems through computational means has been a central topic in computational physics, chemistry, and quantum computing. Mathematically rigorous results that guarantee the efficient computation of ground state and Gibbs state properties typically rely on structural knowledge of these states. For example, the ground state of a 1D gapped local Hamiltonian can be efficiently obtained, as the area-law entanglement scaling ensures that it admits a matrix product state representation [1, 2]. Moreover, Gibbs states can be efficiently prepared on quantum computers when they exhibit exponential decay of correlations and satisfy certain Markov property [3, 4].
Recently, for spin systems, a structural result demonstrated that high temperature Gibbs states are unentangled and can be efficiently prepared by classical computers [5]. More specifically, it was shown that for geometrically local Hamiltonians at a sufficiently high temperature that is independent of the system size, the Gibbs state is a classical distribution over product states and thus exhibits zero entanglement. By efficiently sampling from this distribution, Bakshi et al. [5] extend this structural result into a classical algorithm for sampling and estimating observable values of the Gibbs state. While prior works [6, 7, 8] have shown that high-temperature Gibbs states can be efficiently prepared on quantum computers, the structural result and efficient classical algorithms in [5] suggest that the potential quantum advantage for quantum Gibbs sampling [9, 10, 11] should be investigated in other parameter regimes.
In this work, we investigate analogues of the structural results and classical algorithms for high-temperature Gibbs states for fermionic systems. Fermionic Hamiltonians encompass a broad class of quantum systems of practical interest, including models for interacting electrons such as the electronic structure Hamiltonian [12, 13, 14] and the Fermi-Hubbard model [15, 16]. In contrast to spin systems, where operators acting on different sites commute, fermionic systems are described by Majorana operators, which anti-commute and become non-local when mapped to the qubit setting: using the Jordan–Wigner transformation [17], the $2n$ Majorana operators $\gamma_{1},...,\gamma_{2n}$ can be mapped to tensor products of Pauli operators on $n$-qubit system, with $\gamma_{2j-1}=Z_{1}...Z_{j-1}X_{j},\gamma_{2j}=Z_{1}...Z_{j-1}Y_{j}$. This anti-commutation relation and the non-local nature of Majorana operators induce distinct structural and entanglement properties for fermionic systems, raising the natural question of whether the results obtained for high-temperature Gibbs states in spin systems [5] still hold for fermionic systems. Specifically, are high-temperature fermionic Gibbs states likewise unentangled? Is there an analogous structural result for fermionic systems? If so, can such fermionic Gibbs states also be efficiently prepared using classical algorithms, thereby excluding the potential quantum advantage for these systems?
Main results. We show that at sufficiently high temperatures independent of the system size, the Gibbs state of a geometrically local fermionic Hamiltonian is a classical distribution over a special class of fermionic Gaussian states and can be efficiently sampled. Gaussian states are the ground states and Gibbs states of 2-local fermionic Hamiltonians (free fermions); they are structurally simple, efficiently representable on classical computers, and are often considered the fermionic analogue of product states [18].
More specifically, we prove the following:
Theorem 1.1 (Informal version of Theorem 4.2)
Consider a system of $2n$ Majorana operators on a constant-dimension lattice, and a fermionic Hamiltonian $H=\sum_{a}H_{a}$ where each $H_{a}$ involves only a constant number of Majorana operators on nearby lattice sites. Then, there exists a constant critical temperature such that, above this temperature, the Gibbs state can be written as
$\rho_{\beta}=\frac{e^{-\beta H}}{\text{tr}(e^{-\beta H})}=\sum_{i}q_{i}\rho^{(i)},$ (1)
where $\{q^{(i)}\}_{i}$ is a classical distribution, $\rho^{(i)}$ belongs to a special class of Gaussian states, which can be defined w.r.t a partial matching $M^{(i)}$ of Majorana operators,
$\rho^{(i)}=\frac{1}{2^{n}}\prod_{(k,l)\in M^{(i)}}\left(I+i\gamma_{k}\gamma_{l}\right)$ (2)
[…]
Based on the structural result, we also show an efficient classical algorithm for sampling from the distribution of Gaussian states:
Theorem 1.2 (Informal version of Theorem 5.11)
For the same class of fermionic Hamiltonians $H=\sum_{a}H_{a}$, there exists an efficient classical algorithm to sample from the distribution $q_{i}$ in Theorem 1.1 up to $\epsilon$ precision, and thus the Gibbs states $\rho_{\beta}$ can be efficiently sampled on classical computers.
In the context of the results from [5] for high-temperature Gibbs state in quantum spin systems, one might try to transform the fermionic Hamiltonian to a spin Hamiltonian and then extend their results to the fermionic setting directly. This approach encounters different difficulties based on the different fermionic encodings used. If the Jordan-Wigner transformation is used, then the locality of the Hamiltonian will no longer be preserved except for 1D quantum systems, which violates the locality requirement in [5]. If the locality-preserving encoding of [19] is used, then ancilla spins are needed and these spins need to be in the ground state of an additional Hamiltonian, which violates the high-temperature requirement of [5]. Therefore the results in this work do not follow from such a simple argument.
One may also ask whether the high-temperature fermionic Gibbs states can be represented as a mixture of unentangled states, as done in [5]. Here we present an informal argument, which is independent of the rest of the paper, to show that the fermionic superselection rule precludes such a representation. The fermionic superselection rule forbids superposition between even and odd parity sectors. Mathematically, for a system consisting of modes $\gamma_{1},\gamma_{2},\cdots,\gamma_{2n}$, this means any valid quantum state has to commute with the parity operator $i^{n}\gamma_{1}\gamma_{2}\cdots\gamma_{2n}$. If a fermionic system is decomposed into subsystems $S_{1},S_{2},\cdots,S_{\ell}$ with parity operators $P_{1},P_{2},\cdots,P_{\ell}$ respectively, and a quantum state is unentangled across these subsystems, then it needs to simultaneously commute with parity operators $P_{j}$. If a Gibbs state can be expressed as a mixture of such states, then it also needs to commute with all operators $P_{j}$ by linearity. Since the Gibbs state is proportional to $e^{-\beta H}$, we also need $[H,P_{j}]=0$. This requirement is obviously not satisfied by all Hamiltonians, since a single quadratic term across two subsystems $j,k$ already fails to commute with both $P_{j}$ and $P_{k}$.
Technical highlights. We follow the framework developed for spin systems [5]. Here we highlight the necessary modifications to accommodate the fermionic anti-commutation relations, which lead to an adaptive choice of the matching $M^{(i)}$ in the Gaussian states.
Similar to the qubit case [5], for any sequence of decreasing subsets of Majorana sites, that is, a sequence of subsets $S_{0}\supsetneq S_{1}...\supsetneq S_{2n}$ with $S_{0}=\{1,...,2n\}$ and $S_{2n}=\emptyset$, we define the restricted Hamiltonian $H^{S_{i}}$ to be the Hamiltonian consisting of terms of $H$ whose supports are entirely contained in $S_{i}$. This allows us to decompose the Gibbs states into product of pieces:
$e^{-\beta H}=M_{2n}\dots M_{1}M_{1}^{\dagger}...M_{2n}^{\dagger}\,\text{ for }\,M_{i}:=e^{\beta H^{S_{i}}/2}e^{-\beta H^{S_{i-1}}/2},$ (3)
For easy illustrations, we temporarily assume that $S_{i}=\{i+1,...,2n\}$.
We further show that at high temperatures, each $M_{i}$ is a probability distribution over operators of a simple form. At a high level, we use this simple form to randomly sample a simple operator from $M_{1}M_{1}^{\dagger}$, isolate (pin out) the first site $\gamma_{1}$, and move it to the left of $M_{2n}\dots M_{2}M_{2}^{\dagger}...M_{2n}^{\dagger}$. This procedure leaves us with an operator that acts only on the remaining Majorana operators $[2,3...,2n]$, allowing us to recursively apply similar techniques to isolate subsequent operators.
In the qubit setting, where $S_{i}$ corresponds to the index of qubits and $\gamma_{1}$ corresponds to an operator acting on qubit $1$, it is relatively easy to pin out $\gamma_{1}$ since operators on qubit $1$ commute with the remaining terms $M_{2n}\cdots M_{2}$, which do not act on qubit $1$. However, this approach must be modified in the fermionic setting, where the basic building blocks are Majorana operators that anti-commute rather than commute. For this reason, sites must be pinned in pairs, since operators of the form $\gamma_{k}\gamma_{l}$ do commute with other operators with disjoint supports and even weight. Moreover, two sites can only be pinned out together once the remaining terms in $M_{2n}\dots M_{2n}^{\dagger}$ do not include either site in their support. Pairs of sites are pinned whenever possible, forming a partial matching that is adaptively determined by the random sampling process.
Using the above ideas along with a few additional sampling steps and algebraic manipulations, a distribution of unnormalized fermionic Gaussian states can be obtained that reproduces $e^{-\beta H}$ in expectation. However, it is particularly concerning that their normalization constants could be negative.
To complete the proof, we demonstrate that these normalization constants are actually all nonnegative. The proof of positivity is essential, and it is what separates this approach from a naive sampling of terms from $M_{2n}\dots M_{2n}^{\dagger}$. The proof makes use of a nuance in the algorithm: the $S_{i}$ must be chosen adaptively according to a rule, rather than being set a priori. Likewise, the pairs of pinned indices $(k,l)$ are chosen carefully and adaptively. This latter requirement is a fundamental distinction from the spin system case, in which all the sampled states were product states with respect to the same tensor product decomposition. In the fermionic case, each sampled state is a Gaussian state with respect to a different pairing of Majorana sites, and the state could therefore exhibit more complex correlations between sites than product states in a spin system. And just as in the original proof, the adaptive, algorithmic nature of the decomposition is indispensable in proving an entirely physical result. For a full description of the algorithm, see Section 4.
Open problems. Our results implies that high-temperature Gibbs states of geometrically local fermionic Hamiltonians—and more generally, bounded-degree local fermionic Hamiltonians— is a probability distribution over Gaussian states (mixture of Gaussian states) and can be efficiently simulated on classical computers. It remains to be seen whether our techniques can be extended to rule out quantum advantage in other fermionic settings. In particular, [20, 21] showed that quantum computers can efficiently prepare fermionic Gibbs states at any temperature when the non-quadratic interaction is sufficiently weak. An interesting open question is whether these Gibbs states can also be expressed as mixtures of Gaussian states, thereby enabling efficient classical simulation. Alternatively, it may be shown that in this regime the Gibbs state is not a mixture of Gaussian states, in which case the algorithmic framework developed in [5] would no longer apply. Besides, it is also interesting to investigate whether expressing fermionic Gibbs states as mixture of Gaussian states is related to the performance of quantum Monte Carlo methods in the presence of a sign problem [22, 23, 24]. We can see that the results presented in this work open up numerous directions for future investigation.
Overview. In this paper, we will construct two algorithms, a structural algorithm and a sampling algorithm. The two algorithms rely on different, but both constant, temperature bounds. The structural algorithm does not classically prepare the Gibbs state as a probability distribution of Gaussian states, but its correctness implies the structural result that the Gibbs state lies in the convex hull of Gaussian states.
In particular, the structural algorithm outputs a PSD matrix of the form $\sigma=\frac{\lambda}{\text{tr}(I)}\prod_{(k,l)\in M}(I\pm i\gamma_{k}\gamma_{l})$, where no index is contained in two elements of $M$ (i.e., $M$ is a partial matching of the indices). The matrices $\sigma$ need not be normalized, and can have varying traces $\lambda\geq 0$—they are unnormalized Gaussian states with positive normalization constants. This algorithm, therefore, does not output quantum states at all. However, we will ensure that $\mathbb{E}[\sigma]=e^{-\beta H}$. Despite not preparing normalized quantum states, the existence of this algorithm will algebraically imply that the Gibbs state is a convex combination of Gaussian states.
Indeed, if the algorithm outputs a given $\sigma^{(i)}$ with probability $p_{i}$, then
$e^{-\beta H}=\sum_{i}p_{i}\sigma^{(i)}.$
Each $\sigma^{(i)}=\lambda_{i}\rho^{(i)}$, for a Gaussian state $\rho^{(i)}$ and with $\lambda_{i}=\text{tr}(\sigma^{(i)})$. Defining the nonnegative values
$q_{i}=\frac{p_{i}\lambda_{i}}{\text{tr}(e^{-\beta H})},$
we obtain
$\rho_{\beta}=\sum_{i}q_{i}\rho^{(i)}.$
Taking the trace of both sides of the above equation, we verify that $\sum_{i}q_{i}=1$, proving that $\rho_{\beta}$ is indeed a convex combination of Gaussian states.
After constructing this structural algorithm in Section 4, we will adapt it to an efficient sampling algorithm in Section 5 that can instead efficiently sample $\rho^{(i)}$ with probability $q_{i}$.
Organization. The rest of the paper is organized as follows: In Section 2 we provide the necessary definitions of Majorana operators and the Hamiltonians that we are going to study. In Section 3 we break the Gibbs state into a product of factors, each of which can be efficiently sampled. In Section 4 we present a sampling algorithm over unnormalized fermionic Gaussian states that produces the Gibbs state in expectation value, which proves the structural result that the Gibbs state is a mixture of Gaussian states. In Section 5 we modify the algorithm in the previous section to obtain an efficient algorithm that outputs random normalized fermionic Gaussian states whose expectation value is the Gibbs state.
2 Problem Setup
We use $[2n]$ to denote the set $\{1,2,...,2n\}$. For a subset $S\subseteq[2n]$, we use $|S|$ to denote the number of elements in $S$, and $S^{c}$ to denote the complement set $[2n]\backslash S.$ For two subsets $A,B\subseteq[2n]$, we use the notation $A-B=\{a\in A|\,a\notin B\}$, so that the operation is well-defined even when $B\not\subseteq A$.
A system of $n$ fermion modes can be described by $2n$ Majorana (fermion) operators $\gamma_{j}$, $j\in[2n],$ which satisfy
$\gamma_{j}^{2}=I,\gamma_{j}=\gamma_{j}^{\dagger},tr(\gamma_{j})=0\text{ and }\gamma_{i}\gamma_{j}=-\gamma_{j}\gamma_{i},\forall i\neq j.$
We use Majorana string to refer to a product of an even number of distinct Majorana operators, that is, $G=\alpha\gamma_{i_{1}}\cdots\gamma_{i_{k}}$, with a complex coefficient $\alpha$ and even $k$. We define its support to be $\text{supp}(G):=\{i_{1},...,i_{k}\}$. We define $\mathcal{M}$ to be the set of all $\mathcal{R}$-local Majorana strings with coefficient $i^{\mathcal{R}/2}$, which is used to ensure that the term is Hermitian, that is, $\left(i^{\mathcal{R}/2}\gamma_{i_{1}}...\gamma_{i_{\mathcal{R}}}\right)^{\dagger}=i^{\mathcal{R}/2}\gamma_{i_{1}}...\gamma_{i_{\mathcal{R}}}$. Moreover, we use $\mathcal{M}^{*}$ to denote the set of all Majorana strings with coefficient in {±$\{\pm 1,\pm i\}$ (so either Hermitian or anti-Hermitian Majorana strings).
We say a fermionic Hamiltonian $H$ on $2n$ Majorana operators is of locality $\mathcal{R}$ and degree $d$, if
$H=\sum_{a=1}^{m}\lambda_{a}G_{a},$
where $0\leq\lambda_{a}\leq 1$, $|\text{supp}(G_{a})|$ is even and $\leq\mathcal{R}$, and any local term $G_{a}$ has overlapping support with at most $d$ other terms, i.e. |$|\{b\,|\,\text{supp }(G_{a})\cap\text{supp}(G_{b})\neq\emptyset\}|\leq d$.
Notably, fermionic Gaussian states are the ground states and thermal states of $2$-local fermionic Hamiltonians, and $2$-local fermionic Hamiltonians describe non-interacting fermions and serve as the fermionic counterpart to $1$-local qubit Hamiltonians. Consequently, fermionic Gaussian states can be seen as the natural fermionic analogue of product states.
A special case of Gaussian states can be characterized by a partial matching $M$ of the $2n$ Majorana operators, specified by $m\leq n$ disjoint pairs $(k,l)$ and is defined as
$\rho:=\frac{1}{2^{n}}\prod_{(k,l)\in M}\left(I+i\gamma_{k}\gamma_{l}\right).$ (4)
We use the term unnormalized Gaussian state for an operator to describe a Gaussian state with a positive proportionality constant.
Lastly, we comment on the norms used in the proofs. We denote the operator norm by $\lVert\cdot\rVert_{\infty}$, and the $L_{1}$ distance of probability distributions and the Schatten 1-norm of operators by $\lVert\cdot\rVert_{1}$.
3 Telescoping and Sampling
The algorithm will leverage various telescoping expansions of $e^{-\beta H}$. For any subset $S\subseteq[2n]$, define the restricted Hamiltonian $H^{S}$ as the Hamiltonian consisting of the local terms of $H$ whose supports are entirely contained in $S$. Set $S_{0}=[2n]$ and $S_{2n}=\emptyset$. For a given sequence of properly decreasing subsets $S_{0}\supsetneq\dots\supsetneq S_{2n}$ with $|S_{i}|-|S_{i+1}|=1$, we define
$M_{i}=e^{\beta H^{S_{i}}/2}e^{-\beta H^{S_{i-1}}/2},$ (5)
Note that $S_{0}=[2n]$ and $S_{2n}=\emptyset$, so $H^{S_{2n}}=0$ and $H^{S_{0}}=H$. We can therefore write
$e^{-\beta H}=M_{2n}\dots M_{1}M_{1}^{\dagger}\dots M_{2n}^{\dagger}.$ (6)
The above equation breaks $e^{-\beta H}$ into small pieces. Below, we will provide analysis on the behaviors of this decomposition that are important for our later structural and sampling algorithms. In Lemma 3.1, adapted from Theorem 4.2 in [5], we show that each of $M_{i}$ has a well-behaved Taylor decomposition, where the coefficients decay exponentially in the degree. Corollary 3.3 builds on this by showing that the terms of a fixed degree in the expansion can be efficiently sampled and their support is bounded, which is useful for the later Gibbs sampling algorithm in Section 5.4. Lastly, we will show in Lemma 3.2 that each update step $M_{i}$ in the telescoping expansion can be written as a convex combination of simple, local perturbations of the identity. Lemma 3.2, a consequence of the Taylor decomposition is at the heart of the structural algorithm, because it enables $e^{-\beta H}$ to be decomposed as a convex combination of simple operators.
Lemma 3.1
For any $S_{i}$ and $S_{i-1}$, if $\beta\leq\frac{1}{Cd}$ for some value $C>1$, then the corresponding $M_{i}$ defined in Eq. (5) can be decomposed as $I+\sum_{t>0}\sum_{j}c_{j,t}\Gamma_{j,t}$, where the following properties hold:
  • $\Gamma_{j,t}\in\mathcal{M}^{*}$ is $\mathcal{R}t$-local Majorana string with $\text{supp}(\Gamma_{j,t})\subseteq S_{i-1}$.
  • $c_{j,t}\geq 0$ and $\sum_{j}c_{j,t}\leq C^{-t}.$
Proof.
We can write
$e^{\beta H^{S_{i}}/2}e^{-\beta H^{S_{i-1}}/2}$ $=\left(\sum_{k=0}^{\infty}\frac{(\frac{\beta}{2})^{k}(H^{S_{i}})^{k}}{k!}\right)\left(\sum_{l=0}^{\infty}\frac{(\frac{\beta}{2})^{l}(-H^{S_{i-1}})^{l}}{l!}\right)$ (7)
$=\sum_{t=0}^{\infty}\frac{\beta^{t}}{2^{t}t!}\sum_{k=0}^{t}\frac{(H^{S_{i}})^{k}(-H^{S_{i-1}})^{t-k}t!}{k!(t-k)!}$
$=\sum_{t=0}^{\infty}\frac{\beta^{t}}{2^{t}t!}\underbrace{\sum_{k=0}^{t}{t\choose k}(H^{S_{i}})^{k}(-H^{S_{i-1}})^{t-k}}_{f_{t}(H^{S_{i-1}},H^{S_{i}})}$
Here, the leading term $f_{0}(H^{S_{i-1}},H^{S_{i}})=I$, and the subsequent terms follow the recursive relationship
$f_{t}(H^{S_{i-1}},H^{S_{i}})$ $=H^{S_{i}}f_{t-1}(H^{S_{i-1}},H^{S_{i}})-f_{t-1}(H^{S_{i-1}},H^{S_{i}})H^{S_{i-1}}$ (8)
$=[H^{S_{i}},f_{t-1}(H^{S_{i-1}},H^{S_{i}})]-f_{t-1}(H^{S_{i-1}},H^{S_{i}})(H^{S_{i-1}}-H^{S_{i}})$
The polynomial $p_{t}(H^{S_{i-1}},H^{S_{i}}):=\frac{\beta^{t}}{2^{t}t!}f_{t}(H^{S_{i-1}},H^{S_{i}})$ can be expanded as a nonnegative combinations of Majorana strings in $\mathcal{M}^{*}$. We will do so by using the recursive relations of $f_{t}$, and show by induction that the desired properties are satisfied for this decomposition. In particular, we will show that each $\Gamma_{j,t}$ is a $t$-fold product of operators $G_{a}$. It will follow that $\Gamma_{j,t}$ is $\mathcal{R}t$-local, since each term of the Hamiltonian acts on at most $\mathcal{R}$ sites.
Consider the expansion in Eq. (7). We begin with the base case $t=0$. Since $f_{0}(H^{S_{i-1}},H^{S_{i}})=I$, we define $c_{0,0}=1$, $T_{0,0}=I$ and $c_{j,0}=0,T_{j,0}=0$ $\forall j\geq 1$, given that the identity operator trivially satisfies the locality requirement with coefficient $1$. For the inductive case, assume for some $t\geq 0$, we have
$p_{t}(H^{S_{i-1}},H^{S_{i}})=\sum_{j}c_{j,t}\Gamma_{j,t}$
with the desired properties. That is, $\Gamma_{j,t}$ is a $t$-fold product of operators $G_{a}$, so $\mathcal{R}t$-local, $c_{j,t}\geq 0$, and $\sum_{j}c_{j,t}\leq C^{-t}.$ Directly applying the recursive relationship in Eq. (8), we have that
$p_{t+1}(H^{S_{i-1}},H^{S_{i}})=\frac{\beta}{2(t+1)}\sum_{j}c_{j,t}\left([H^{S_{i}},\Gamma_{j,t}]-\Gamma_{j,t}\left(H^{S_{i-1}}-H^{S_{i}}\right)\right).$ (9)
In the following, we will simplify Eq. (9) by observing that the commutator $[H^{S_{i}},\Gamma_{j,t}]$ is non-zero only on the local terms where $H^{S_{i}}$ and $\Gamma_{j,t}$ overlap. Define $T_{1,j}$ as the set of local terms of $H^{S_{i}}$ whose support overlaps with $\Gamma_{j,t}$:
$T_{1,j}:=\{b:\text{supp}(G_{b})\subseteq S_{i}\text{ and }[G_{b},\Gamma_{j,t}]\neq 0\}.$
Note that
$[H^{S_{i}},\Gamma_{j,t}]$ $=\sum_{a\in T_{1,j}}(2\lambda_{a})\frac{[G_{a},\Gamma_{j,t}]}{2}$
$=\sum_{a\in T_{1,j}}(2\lambda_{a})G_{a}\Gamma_{j,t},$
since the expression $\frac{[G_{a},\Gamma_{j,t}]}{2}$ either equals zero, when the two terms commute, or equal to $G_{a}\Gamma_{j,t}$ when they anti-commute.
Moreover, recall that by definition $S_{i-1}$ has just one more element than $S_{i}$. Then, the local $G_{b}$ in $H^{S_{i-1}}-H^{S_{i}}$ can be characterized by the following set
$T_{2}:=\{b:v\in\text{supp}(G_{b})\text{ for }\{v\}=S_{i-1}-S_{i}\text{ and }G_{b}\in H^{S_{i-1}}\}.$
We have therefore obtained a decomposition
$p_{t+1}(H^{S_{i-1}},H^{S_{i}})=\frac{\beta}{2(t+1)}\left(\sum_{a\in T_{1,j},j}(2\lambda_{a}c_{j,t})(G_{a}\Gamma_{j,t})+\sum_{j,a\in T_{2}}(\lambda_{a}c_{j,t})(-\Gamma_{j,t}G_{a})\right).$ (10)
Further note that by induction, every Majorana string in Eq. (10) is a $(t+1)$-fold product of terms $G_{a}$ and is thus at most $\mathcal{R}(t+1)$-local. Moreover, note that $\text{supp}(\Gamma_{j,t})\subseteq S_{i-1}$ since both $H^{S_{i}}$ and $H^{S_{i-1}}$ are supported on $S_{i-1}.$
Finally, we prove that for this decomposition, $\sum_{j^{\prime}}c_{j^{\prime},(t+1)}\leq C^{-(t+1)}$. To do this, we will bound the sizes of $T_{1,j},\forall j$ and the size of $T_{2}$. First, note that for any $\Gamma_{j,t}$, which is a $t$-fold product of Majorana strings $G_{a}$, there are at most $td$ operators $G_{a}$ in the Hamiltonian that have a nonzero commutator with $\Gamma_{j,t}$. It follows that $|T_{1,j}|\leq td$. On the other hand, each $a\in T_{2}$ corresponds to a $G_{a}$ whose support includes the one index $v\in S_{i-1}-S_{i}$. All of these operators have intersecting support, which means $|T_{2}|\leq d+1$. We may now write
$\sum_{j^{\prime}}c_{j^{\prime},(t+1)}$ $=\frac{\beta}{2(t+1)}\sum_{a\in T_{1,j}}\sum_{j}2\lambda_{a}c_{j,t}+\frac{\beta}{2(t+1)}\sum_{a\in T_{2}}\sum_{j}\lambda_{a}c_{j,t}$ (11)
$\leq\frac{\beta}{2(t+1)}\sum_{j}c_{j,t}(2td+d+1)$
$\leq\frac{\beta}{2(t+1)}\sum_{j}2(t+1)c_{j,t}d$
$\leq\frac{1}{C}\sum_{j}c_{j,t}.$
where the first inequality uses that $d\geq 1$ and the last inequality uses that $\beta\leq\frac{1}{Cd}$. Combining with the inductive hypothesis that $\sum_{j}c_{j,t}\leq C^{-t}$, we arrive at the desired bound.
Lemma 3.2
Assume that $\beta\leq\frac{1}{2N\mathcal{R}d}$ for some $N>0$, then each operator
$M_{i}=e^{\beta H^{S_{i}}/2}e^{-\beta H^{S_{i-1}}/2}$
and correspondingly $M_{i}^{\dagger}$ is a probability distribution over operators of the form $I+(N\mathcal{R})^{-t_{k}}\Gamma_{k}$, where for any $k$, $\Gamma_{k}\in\mathcal{M}^{*}\cup\{0\}$ is $t_{k}\mathcal{R}$-local and is supported on $S_{i-1}$.
Proof.
We know from Lemma 3.1 that $M_{i}$ can be written as $I+\sum_{t>0}\sum_{j}c_{j,t}\Gamma_{j,t}$. Then we can rewrite $M_{i}$ as
$M_{i}$ $=\left(1-\sum_{t>0}\sum_{j}c_{j,t}(N\mathcal{R})^{t}\right)I+\sum_{t>0}\sum_{j}\left(c_{j,t}(N\mathcal{R})^{t}\right)\left(I+(N\mathcal{R})^{-t}\Gamma_{j,t})\right).$ (12)
Note that by Lemma 3.1 we know that $\sum_{j}c_{j,t}\leq(2N\mathcal{R})^{-t}$, $c_{j,t}\geq 0$, thus in Eq. (12) the coefficients of $I$ and $\left(I+(N\mathcal{R})^{-t}\Gamma_{j,t})\right)$ are non-negative and sum to one. Therefore, Eq. (12) shows that $M_{i}$ is a probability distribution over operators $I+(N\mathcal{R})^{-t_{k}}\Gamma_{k}$, where to ease notation we treat $I=I+(N\mathcal{R})^{0}0$ as the term $k=0$, and we relabel the indexes $(j,t)$ for $t>0$ arbitrarily as $k=1,2,...$.
Moreover, by Lemma 3.1 we know that for $t>0$, $\Gamma_{j,t}$ is $\mathcal{R}t$-local and is supported on $S_{i-1}$, thus the desired properties hold. The desired properties trivially holds for the term $I=I+(N\mathcal{R})^{0}0$. Thus we complete the proof, and the same argumentation applies almost identically to $M_{i}^{\dagger}$.
The above lemma will be used with $N=24$ to prove the structural result, and with $N=25$ to prove the sampling result. This is where the bounds on inverse temperature $\beta$ arise from.
Lastly, we record a corollary which will be useful for the later Gibbs sampling algorithm in Section 5.
Corollary 3.3
For any fixed $t$ and $\beta\leq\frac{1}{Cd}$, the values $(c_{j,t},\Gamma_{j,t})$ can be sampled in time $\text{poly}(t,n)$ according to a distribution $\mu$ such that $\mu_{j,t}\geq c_{j,t}C^{t}$, and which has support of size at most $(t+1)!(d+1)^{t}$, where $d$ is the degree of the interaction graph of $H$.
Proof.
The distribution $\mu$ will be sampled by a repeated application of Equation 10. We will prove this result by induction. For $t=0$, the only term in the decomposition is the identity operator $I$ with coefficient $c_{0,0}=1$, corresponding to $\Gamma_{0,0}=I.$ We define $\mu_{0,0}=1.$ Thus,
$\mu_{0,0}=c_{0,0}\geq c_{0,0}C^{0},$
and the support size is $1$. The base case holds.
Suppose that for some $t\geq 0,$ there exists a distribution $\mu$ over pairs $(c_{j,t},\Gamma_{j,t})$ that can be sampled in $poly(t,n)$ time such that $\mu_{j,t}\geq c_{j,t}C^{t}$ for all $j$ and the support size is at most $(t+1)!(d+1)^{t}$. We will construct such a distribution $\mu^{\prime}$ for degree $t+1.$ Given a sampled pair $(c_{j,t},\Gamma_{j,t})$ from $\mu$ at degree $t$, we extend it to a term $(c_{j^{\prime},t+1},\Gamma_{j^{\prime},t+1})$ at degree $t+1$ by selecting some $a\in T_{1,j}$, or some $a\in T_{2}$. This yields
$c_{j^{\prime},t+1}=\frac{\beta}{t+1}\lambda_{a}c_{j,t}$
or
$c_{j^{\prime},t+1}=\frac{\beta}{2(t+1)}\lambda_{a}c_{j,t}$
respectively, with $\Gamma_{j^{\prime},t}$ accordingly set to either $G_{a}\Gamma_{j,t}$ or $-\Gamma_{j,t}G_{a}$. Now, $\mu_{j^{\prime},t+1}$ is the distribution obtained by first sampling $(c_{j,t},\Gamma_{j,t})$ from $\mu$, and then sampling a random value uniformly from $T_{1,j}$ with probability $P=\frac{2td}{2td+(d+1)}$ or from $T_{2}$ with probability $1-P=\frac{d+1}{2td+(d+1)}$. Note that $P\geq\frac{td}{(t+1)d}$ and $1-P\geq\frac{d+1}{2(t+1)d}$. Now, if $(j,t+1)$ corresponds to a value chosen from $T_{1,j}$, we have
$\mu_{j^{\prime},t+1}$ $\geq\frac{P}{td}\mu_{j,t}\geq\frac{\beta C}{t+1}\mu_{j,t}\geq C^{t+1}\frac{\beta}{t+1}c_{j,t}\geq C^{t+1}c_{j^{\prime},t+1},$
where first inequality comes from $|T_{1,j}|\leq dt$, the second inequality uses $1/d\geq\beta C$, the third inequality comes from the inductive hypothesis, and the last inequality follows from the fact that $0\leq\lambda_{a}\leq 1$.
On the other hand, if $(c_{j^{\prime},t+1},\Gamma_{j^{\prime},t+1})$ corresponds to a value chosen from $T_{2}$, then using the bound $|T_{2}|\leq d+1$, we obtain
$\mu_{j^{\prime},t+1}$ $\geq\frac{1-P}{d+1}\mu_{j,t}\geq\frac{1}{2(t+1)d}\mu_{j,t}\geq\frac{\beta C}{2(t+1)}\mu_{j,t}\geq C^{t+1}\frac{\beta}{2(t+1)}c_{j,t}\geq C^{t+1}c_{j^{\prime},t+1},$
as desired.
Note that every step in the above runs in time that is polynomial in $n$ except potentially the uniform sampling of $T_{1,j}$ and $T_{2}$. However, this step can also be done efficiently. Specifically, the set $T_{2}$ contains at most $d+1$ elements, which can be easily enumerated from an encoding of the terms in $H$. The set $T_{1,j}$ contains at most $td$ elements, corresponding to local Hamiltonian terms that overlap with the support of $\Gamma_{j,t}$. Each element can be identified in time $O(t)$ by checking if it anti-commutes with $\Gamma_{j,t}$. Once these elements are obtained, sampling uniformly from them is efficient.
We can also analyze the support of the sampling distribution. For each $t^{\prime}<t$, there is a choice of one of at most $(t^{\prime}+1)d+1\leq(t^{\prime}+1)(d+1)$ indices. For $t$ itself, it follows that there are at most $(t+1)!(d+1)^{t}$ elements in the support. ∎
4 Structural Algorithm
In this section, we show that above a high temperature threhold that is constant of system size, the Gibbs state of fermionic local Hamiltonians is exactly a probability distribution over Gaussian states. We first provide an overview in Section 4.1, then introduce the adaptive algorithm used to prove the structural result in Section 4.2. The formal proof of correctness of the algorithm is presented in Section 4.3.
4.1 Overview
The algorithm below stochastically prepares an unnormalized fermionic Gaussian state $\sigma$, such that the output distribution has expectation $\mathbb{E}[\sigma]=e^{-\beta H}$. To do so, it iteratively expands $e^{-\beta H}$ using a stochastically chosen telescoping set of operators $M_{j}$, as in Section 3. In particular, at each step $j$, the algorithm chooses a site $i$ to remove from $S_{j}$, and defines $M_{j+1}\coloneq e^{\beta H^{S_{j+1}}}e^{-\beta H^{S_{j}}}$ with $S_{j+1}=S_{j}-\{i\}$.
Using this telescoping product, the algorithm iteratively prepares operators $\Gamma_{j}\in\mathcal{M},\sigma_{j}$, and a scalar $\alpha_{j}\geq 0$ at each step $j$, that satisfy the rule
$\mathbb{E}\left[M_{1}^{-1}\dots M_{j}^{-1}(I+\alpha_{j}\Gamma_{j})\sigma_{j}(M_{j}^{{\dagger}})^{-1}\dots(M_{1}^{{\dagger}})^{-1}\right]=I,$ (13)
taking an expectation over the randomness of the algorithm. It is always true that $\sigma_{j}$ is an unnormalized separable state. Moreover, the variables are initialized with $\sigma_{0}=I$, $\Gamma_{0}=0$, and $\alpha_{0}=0$, while it is importantly ensured that $\Gamma_{2n}=0$. This latter condition means that by the final step $j=2n$, we have $M_{2n}M_{2n-1}\cdots M_{1}=e^{-\beta H/2}$, and the formula reads
$\mathbb{E}\left[e^{\beta H/2}\sigma_{2n}e^{\beta H/2}\right]=I,$
from which it follows that $\mathbb{E}[\sigma_{2n}]=e^{-\beta H}$.
How does this decomposition work? The algorithm’s goal is to build up a distribution of separable states, $\sigma_{2n}$, that on expectation is equal to the Gibbs state. It does this by iteratively constructing operators $(I+\alpha_{j}\Gamma_{j})\sigma_{j}$, where $I+\alpha_{j}\Gamma_{j}$ is supported on sites that have yet to be extracted, and $\sigma_{j}$ is a separable state on the sites that have already been extracted. We introduce the terminology of “pinnable” and “pinned” sites. A site $i$ is termed “pinnable” when $i\notin S_{j}^{c}$, and it is termed “pinned” when $i\notin S_{j}^{c}$ and $i\notin\Gamma_{j}$, so it has already been extracted from $\Gamma_{j}$.
It is crucial that any site extracted to $\sigma_{j}$ have support contained in $S_{j}^{c}$, because this ensures it has disjoint support with every subsequent $M_{k}$, $k>j$. This is why only pinnable sites should be extracted to $\sigma_{j}$. When a site has been extracted to $\sigma_{j}$, or if it is naturally absent from the support of $\Gamma_{j}$ and not contained in $S_{j}$, it is “pinned”. In either case, it is fully extracted, and is ensured to never reappear in subsequent $M_{k}$ operators.
Let us see an example. Since $\sigma_{0}=I$ and $\Gamma_{0}=0$, our algorithm begins with the operator $(I+\alpha_{0}\Gamma_{0})\sigma_{0}=I$. Then, in the first step, it conjugates this operator, in expectation value, by a well-chosen $M_{1}$ (for some site $i$), yielding
$M_{1}(I+\alpha_{0}\Gamma_{0})\sigma_{0}M_{1}^{\dagger}.$
We may bring $\sigma_{0}$ to the side, since it commutes with $M_{1}$. Notice that by the sampling argument in Lemma 3.2, we can also sample Majorana operators $\beta_{1}\Lambda_{1}$ and $\beta_{2}\Lambda_{2}$ from $M_{1}$ and $M_{1}^{\dagger}$ here, to obtain an expression
$I+\alpha_{1}\Gamma_{1}=(I+\beta_{1}\Lambda_{1})(I+\alpha_{0}\Gamma_{0})(I+\beta_{2}\Lambda_{2})\sigma_{0}$ (14)
for which $\mathbb{E}[M_{1}^{-1}(I+\alpha_{1}\Gamma_{1})\sigma_{1}(M_{1}^{\dagger})^{-1}]=I$, where $\sigma_{1}=\sigma_{0}$. To simplify this expression further, however, we must perform additional sampling. Our expression can be foiled out as
$c_{1}(I+c_{1}^{-1}\beta_{1}\Lambda_{1})\sigma_{0}+\dots+c_{6}(I+(c_{6}^{-1}\beta_{1}\alpha_{0}\beta_{2})(\Lambda_{1}\Gamma_{j}\Lambda_{2})\sigma_{0})$
[…]
The first step of the algorithm is almost done: however, we observe there is now one new pinnable site, namely site $i$. Because site $i$ is pinnable, if we could remove it from the support of $\Gamma_{1}$ and extract it to $\sigma_{1}$ somehow, this would permanently remove site $i$ from future $\Gamma_{j}$ operators. After all, $i\notin\text{supp}(M_{k})$ for $k\geq 2$. And indeed, if there is pair of pinnable unpinned sites that are both present in $\Gamma_{1}$, i.e. if $\Gamma_{1}=\Gamma_{1}^{-}(i\gamma_{i}\gamma_{k})$ for some Majorana operators $\gamma_{i},\gamma_{k}$ with $k\notin S_{1}$, then we may write
$(I+\alpha_{1}\Gamma_{1})\sigma_{0}=\frac{1}{2}(I+\alpha_{1}\Gamma_{1}^{-})(I+i\gamma_{i}\gamma_{k})\sigma_{0}+\frac{1}{2}(I-\alpha_{1}\Gamma_{1}^{-})(I-i\gamma_{i}\gamma_{k})\sigma_{0}.$ (15)
Sampling one of these two terms, setting $\Gamma_{1}$ to be ±$\pm\Gamma_{1}$ and $\sigma_{1}=(I\pm i\gamma_{i}\gamma_{k})\sigma_{0}$ consistently, we have extracted site $i$ and completed one step of the algorithm! Since eventually every site must be pinned, by the end of the algorithm $\Gamma_{2n-1}$ will have trivial support, i.e. will be a scalar multiple of the identity. At this point, $(I+\alpha_{2n-1})\Gamma_{2n-1}$ is a scalar, so defining $\sigma_{2n}\coloneq(I+\alpha_{2n-1})\Gamma_{2n-1}\sigma_{2n-1}$ yields an unnormalized Gaussian state, as desired.
The biggest impediment to the strategy outlined above is that it must be ensured that the constant term absorbed into $\sigma_{2n-1}$ is nonnegative. The algorithm, as currently described, could end with $(I+\alpha_{2n-1}\Gamma_{2n-1})\sigma_{2n-1}$ such that $\alpha_{2n}>1$, for which it is possible that the scalar $I+\alpha_{2n-1}\Gamma_{2n-1}<0$, and therefore the operator $\sigma_{2n}$ is not PSD. This is an essential detail—it is the reason why the specific bounds on the sampled operators in Lemma 3.2 matter, and precisely why this approach only works for sufficiently high-temperature Gibbs states.
To remedy this problem, the algorithm must control the growth of $\alpha_{j}$ in any way possible. The choice of the sampling parameters $c_{i}$ is helpful in preventing the values $\alpha_{j}$ from growing past 1. The algorithm is also altered to make it possible to periodically reset $\Gamma_{j}$ and $\alpha_{j}$ to zero. Note that whenever $\Gamma_{j}$ has empty support, $I+\alpha_{j}\Gamma_{j}$ is just a scaling factor. This factor can therefore be brought into $\sigma_{j}$ whenever $\Gamma_{j}$ has empty support, not just in the last step. By redefining $\sigma_{j}\leftarrow(I+\alpha_{j}\Gamma_{j})\sigma_{j}$ and $(\Gamma_{j},\alpha_{j})\leftarrow(0,0)$, $(I+\alpha_{j}\Gamma_{j})\sigma_{j}$ is left unchanged. As long as $\alpha_{j}$ is at most 1 whenever this step is taken, $\sigma_{j}$ remains PSD, and $\alpha_{j}$ is now reset to 0, slowing its growth. The algorithm will always choose to pin a site from $S_{j}$ that is in the support of $\Gamma_{j}$. This has the effect of driving the support closer to empty. It is precisely for this reason that the algorithm must adaptively choose which site to pin in order to succeed. A proof of this result is given in Lemma 4.8.
4.2 Algorithmic Description
Now, here is the algorithmic description.
Algorithm 4.1 (Structural Algorithm)
Input: Fermionic Hamiltonian $H=\sum_{a}\lambda_{a}G_{a}$ with locality $\mathcal{R}$ and degree $d$; inverse temperature parameter $\beta$ such that $\beta\leq\frac{1}{2N\mathcal{R}}$ for $N>0$.
Output: An unnormalized Gaussian state $\sigma$.
  1. 1.
    Initialize $\sigma_{0}=I,\Gamma_{0}=0,\alpha_{0}=0,S_{0}=[2n]$.
  2. 2.
    For $j$ in $10,\dots,2n-$ 1:
    1. (a)
      If $\text{supp}(\Gamma_{j})\cap S_{j}$ is nonempty, set $i$ as the minimal index in $\text{supp}(\Gamma_{j})\cap S_{j}$. Otherwise, let $i$ be the minimal index in $S_{j}$. Define $S_{j+1}=S_{j}-\{i\}$. Define
      $M_{j+1}=e^{\beta H^{S_{j+1}}/2}e^{-\beta H^{S_{j}}/2}.$
    2. (b)
      Sample $I+\beta_{1}\Lambda_{1}$ for $\beta_{1}=(N\mathcal{R})^{-t_{1}}$ from the probability distribution in Lemma 3.2 w.r.t $M_{j+1}$, so that [^1] For simplicity, with a slight abuse of notation, we denote this sampled $t_{k}$ as $t_{1}$ and write $\beta_{1}=(N\mathcal{R})^{-t_{1}}$.
      $\mathbb{E}\left[I+\beta_{1}\Lambda_{1}\right]=M_{j+1}=e^{\beta H^{S_{j+1}}/2}e^{-\beta H^{S_{j}}/2}.$
    3. (c)
      Sample $I+\beta_{2}\Lambda_{2}$ for $\beta_{2}=(N\mathcal{R})^{-t_{2}}$ independently from the same distribution, so that
      $\mathbb{E}\left[I+\beta_{2}\Lambda_{2}\right]=M_{j+1}^{\dagger}.$
    4. (d)
      Set $\sigma_{j+1}\leftarrow\sigma_{j}$ and $(\Gamma_{j+1},\alpha_{j+1})$ as:
      1. i.
        $\bigl{(}\Gamma_{j},(1-\frac{1}{\mathcal{R}})^{-1}\alpha_{j}\bigl{)}$ with probability $1-\frac{1}{\mathcal{R}}$
      2. ii.
        $\left(\Lambda_{1},(6\mathcal{R})\beta_{1}\right)$ with probability $\frac{1}{6\mathcal{R}}$
      3. iii.
        $\left(\Lambda_{2},(6\mathcal{R})\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
      4. iv.
        $\left(\Lambda_{1}\Gamma_{j},(6\mathcal{R})\beta_{1}\alpha_{j}\right)$ with probability $\frac{1}{6\mathcal{R}}$
      5. v.
        $\left(\Gamma_{j}\Lambda_{2},(6\mathcal{R})\alpha_{j}\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
      6. vi.
        $\left(\Lambda_{1}\Lambda_{2},(6\mathcal{R})\beta_{1}\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
      7. vii.
        $\left(\Lambda_{1}\Gamma_{j}\Lambda_{2},(6\mathcal{R})\beta_{1}\alpha_{j}\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    5. (e)
      If $\Gamma_{j+1}$ is not Hermitian:
      • $\Gamma_{j+1}\leftarrow 0,\alpha_{j+1}\leftarrow 0$.
    6. (f)
      If there are two indices $k,l\in S_{j}^{c}$ such that $k,l\in\text{supp}\left(\Gamma_{j+1}\right)$, i.e. $\Gamma_{j+1}=\Gamma^{-}_{j+1}\cdot(i\gamma_{k}\gamma_{l})$ for some operator $\Gamma^{-}_{j+1}$ whose support is $\text{supp}(\Gamma_{j+1})\backslash\{k,l\}$,:
      1. i.
        $\Gamma_{j+1}\leftarrow\Gamma^{-}_{j+1}$, $\sigma_{j+1}\leftarrow(I+i\gamma_{k}\gamma_{l})\cdot\sigma_{j+1}$ with probability $\frac{1}{2}$
      2. ii.
        $\Gamma_{j+1}\leftarrow-\Gamma^{-}_{j+1}$, $\sigma_{j+1}\leftarrow(I-i\gamma_{k}\gamma_{l})\cdot\sigma_{j+1}$ with probability $\frac{1}{2}$
    7. (g)
      If $\Gamma_{j+1}\in\{\pm I\}$:
      • $\sigma_{j+1}\leftarrow(I+\alpha_{j+1}\Gamma_{j+1})\sigma_{j+1}$, $\Gamma_{j+1}\leftarrow 0$, and $\alpha_{j+1}\leftarrow 0$.
  3. 3.
    Set $\sigma\leftarrow\sigma_{2n}$.
The output of the above algorithm is summarized as follows:
Theorem 4.2
Consider a fermionic Hamiltonian $H=\sum_{a}\lambda_{a}G_{a}$ with locality $\mathcal{R}$ and degree $d$, and an inverse temperature parameter $\beta$ such that $\beta\leq\frac{1}{2N\mathcal{R}d}$ for $N\geq 24$, the random variable $\sigma$ generated by Algorithm 4.1 is always an unnormalized Gaussian state. Moreover, taking into account the randomness in the algorithm, we have
$\mathbb{E}[\sigma]=e^{-\beta H}.$
We prove Theorem 4.2 in Section 4.3.
4.3 Analysis
4.3.1 Notation
First, we establish notation and definitions. Let $\Gamma_{j+1}^{(x)}$ denotes the value of $\Gamma_{j+1}$ at the end of step $(x)$ of the $j$th iteration of the algorithm, for $x\in\{a,b,...,g\}$. Meanwhile, $\Gamma_{j}$ simply refers to the final value it is set to, i.e. $\Gamma_{j}^{(g)}$. Similar notational conventions are adopted for $\sigma_{j}$ and $\alpha_{j}$. Moreover, denote $T_{j}=\text{supp}(\Gamma_{j})$ and the pinned sites $P_{j}=\text{supp}(\sigma_{j})$. Recall that $S_{0}=[2n]$ and the sets $S_{1},...,S_{2n-1}$ are random variables defined by the Algorithm 4.1.
We will also define a filtration $\{\mathcal{F}_{j}^{(x)}\}$. Each $\mathcal{F}_{j}^{(x)}$ is a $\sigma$-algebra that represents all probabilistic events that can be expressed in terms of the randomness in the algorithm until the end of step $(x)$ of the $j$th iteration. In other words, $\mathcal{F}_{j}^{(x)}$ represents the information about the outcomes of random choices up until this point. In particular, we abbreviate $\mathcal{F}_{j}=\mathcal{F}_{j}^{(g)}$, the randomness until iteration $j-1$ ends. We also extend our definition of $\mathcal{F}_{j}$ to $j=0$ by defining $\mathcal{F}_{0}$ to be the trivial $\sigma$-algebra, since at the beginning of the algorithm no stochastic choices have occurred yet. We may then define the conditional expectation $\mathbb{E}\left[X|\mathcal{F}_{j}^{(x)}\right]$ to be the average of $X$ over any uncertainty not contained in $\mathcal{F}_{j}^{(x)}$. It may be thought of as the best guess for $X$, conditioned on the knowledge of whether any event in $\mathcal{F}_{j}^{(x)}$ occurred or did not occur. For instance,
$\mathbb{E}\left[X|\mathcal{F}_{0}\right]=\mathbb{E}\left[X\right],$
since the trivial $\sigma$-algebra $\mathcal{F}_{0}$ contains no information, so the best guess is simply the normal expectation. The other extreme is that
$\mathbb{E}\left[X|\mathcal{F}\right]=X$
whenever the random variable $X$ is perfectly described by the information in $\mathcal{F}$. In this case, $X$ is called $\mathcal{F}$-measurable. For instance, if $X$ is defined deterministically by stochastic choices made up to step $x$ of iteration $j$, then $X$ is $\mathcal{F}_{j}^{(x)}$-measurable. One convenient property that we will make use of is that if $X$ is $\mathcal{F}$-measurable, then
$\mathbb{E}[XY|\mathcal{F}]=X\mathbb{E}[Y|\mathcal{F}].$
In other words $\mathbb{E}[\cdot|\mathcal{F}]$ is linear with respect to $\mathcal{F}$-measurable random variables.
If step $x_{1}$ of iteration $j_{1}$ occurs before step $x_{2}$ of iteration $j_{2}$, then $\mathcal{F}_{j_{1}}^{(x_{1})}\subseteq\mathcal{F}_{j_{2}}^{(x_{2})}$. In general, for any two $\sigma$-algebras if $\mathcal{G}_{1}\subseteq\mathcal{G}_{2}$, then for some random variable $X$,
$\mathbb{E}\left[\mathbb{E}\left[X|\mathcal{G}_{2}\right]|\mathcal{G}_{1}\right]=\mathbb{E}\left[X|\mathcal{G}_{1}\right].$
This is the law of total expectation. In our setting, it expresses that averaging over the randomness beginning after step $(j_{2},x_{2})$, then averaging over the randomness beginning after $(j_{1},x_{1})$, is equivalent to simply averaging over the randomness beginning after $(j_{1},x_{1})$. Each of the above properties will prove useful in the rigorous probabilistic analysis of the algorithm.
4.3.2 Proof of Correctness
To aid our proof, we will show two statements concerning the relationship between the sets $S_{j},T_{j},$ and $P_{j}$ in the following Lemma 4.3. Firstly, the pinned sites, $P_{j}$, are disjoint from $S_{j}\cup T_{j}$, which are the remaining sites that could be in the support of any $M_{i}$ for $i>j$ or in the support of $\Gamma_{j}$. Secondly, $|T_{j}-S_{j}|\leq 1$, which has the interpretation that the support of $\Gamma_{j}$ can only include one element outside of $S_{j}$. In other words, there is only one site $i$ that is pinnable, i.e. $i\in S_{j}^{c}$, but unpinned, i.e. $i\notin P_{j}$. Here, we are using the notational convention that if $A$ and $B$ are sets, then $A-B=\{a\in A|\,a\notin B\}$, so that the operation is well-defined even when $B\not\subseteq A$.
Lemma 4.3 indicates that as $S_{j}$ becomes smaller and smaller, the support of $\Gamma_{j}$ eventually dwindles as well.
Lemma 4.3
For any $j\in 0,\dots,2n$, $(S_{j}\cup T_{j})\cap P_{j}=\emptyset$, and $|T_{j}-S_{j}|\leq 1$.
Proof.
We can prove $(S_{j}\cup T_{j})\cap P_{j}=\emptyset$ by induction. Firstly, for $j=0$, $P_{0}=\emptyset$ since $\sigma_{0}=I$.
Now, note that
$S_{j+1}\cup T_{j+1}\subseteq S_{j}\cup T_{j+1}\subseteq S_{j}\cup T_{j}.$ (16)
The first inclusion is true because $S_{j+1}\subseteq S_{j}$. For the second inclusion, the support of $\Gamma_{j+1}$ is contained in the support of $\Gamma_{j+1}^{(d)}$, which is some product of $\Lambda_{1},\Lambda_{2},$ and $\Gamma_{j}$, and therefore the support is contained in the union of $\text{supp}(\Gamma_{j})$ (which is $T_{j}$ by definition), $\text{supp}(\Lambda_{1})\subseteq S_{j}$, and $\text{supp}(\Lambda_{2})\subseteq S_{j}$, where the last two inclusions come from Lemma 3.2. Therefore, we also show that
$T_{j+1}\subseteq\text{supp}\left(\Gamma_{j+1}^{(d)}\right)\subseteq S_{j}\cup T_{j},$ (17)
so that $S_{j}\cup T_{j+1}\subseteq S_{j}\cup T_{j}$.
Decomposing $P_{j+1}=P_{j}\cup(P_{j+1}-P_{j})$, we may rewrite
$(S_{j+1}\cup T_{j+1})\cap P_{j+1}$ $\subseteq\left((S_{j}\cup T_{j+1})\cap P_{j}\right)\cup\left((S_{j}\cup T_{j+1})\cap(P_{j+1}-P_{j})\right)$
$\subseteq\left((S_{j}\cup T_{j})\cap P_{j}\right)\cup\left((S_{j}\cup T_{j+1})\cap(P_{j+1}-P_{j})\right)$
$=(S_{j}\cup T_{j+1})\cap(P_{j+1}-P_{j})$
where the second $\subseteq$ follows from Eq. (16), and the last equality holds by the inductive hypothesis. Furthermore, note that the elements that can be added from $P_{j}$ to $P_{j+1}$ (in step $(f)$) are indices $k,l$ such that $k,l\notin S_{j}$, and by construction $k,l\notin T_{j+1}$, since $\gamma_{k}$ and $\gamma_{l}$ are removed from the support of $\Gamma_{j+1}$. It follows that $(S_{j}\cup T_{j+1})\cap(P_{j+1}-P_{j})=\emptyset$, which completes the inductive hypothesis.
Now, we would like to prove that $|T_{j}-S_{j}|\leq 1$, again by induction. Since $S_{0}=[2n]$, the base case follows immediately. Now, we assume the induction hypothesis. In Eq. (17) we showed that $T_{j+1}\subseteq\text{supp}\left(\Gamma_{j+1}^{(d)}\right)\subseteq S_{j}\cup T_{j}$, so $T_{j+1}-S_{j}\subseteq\text{supp}\left(\Gamma_{j+1}^{(d)}\right)-S_{j}\subseteq T_{j}-S_{j}$. It follows that
$T_{j+1}-S_{j+1}$ $\subseteq(T_{j+1}-S_{j})\cup(S_{j}-S_{j+1})$
$\subseteq(T_{j}-S_{j})\cup(S_{j}-S_{j+1})$ (18)
The set $S_{j}-S_{j+1}$ consists of one element, say $S_{j}-S_{j+1}=\{i\}$. If $|T_{j}-S_{j}|=0$, then $T_{j+1}-S_{j+1}\subseteq S_{j}-S_{j+1}=\{i\}$, so $|T_{j+1}-S_{j+1}|\leq 1$. On the other hand, if $|T_{j}-S_{j}|=1$, then potentially the largest value for $|T_{j+1}-S_{j+1}|$ is two. However, in this case, by $T_{j+1}\subseteq\text{supp}(\Gamma_{j+1}^{(d)})$, it would also be true that $\left|\text{supp}\left(\Gamma_{j+1}^{(d)}\right)-S_{j+1}\right|\geq 2$, and therefore the cardinality is precisely 2 since $\text{supp}\left(\Gamma_{j+1}^{(d)}\right)-S_{j+1}\subseteq(S_{j}\cup T_{j})-S_{j+1}\subseteq(T_{j}-S_{j})\cup\{i\}$. This implies that there are precisely two indices in the support of $\Gamma_{j+1}^{(d)}$ that are in $S_{j+1}^{c}$. However, these two indices would then have been pinned immediately in step $(f)$, so that $\left|T_{j+1}-S_{j+1}\right|=0$. This completes the proof. ∎
Several corollaries follow easily from this lemma.
Corollary 4.4
Each $\sigma_{j}$, and in particular the output $\sigma=\sigma_{2n}$ of the algorithm, is an unnormalized Gaussian state.
Proof.
Whenever step $(f)$ is run and a term of the form $I\pm i\gamma_{k}\gamma_{l}$ is factored into $\sigma_{j+1}$, we have that $k,l\in\text{supp}\left(\Gamma_{j+1}^{(d)}\right)\subseteq T_{j}\cup S_{j}$, implying that $k,l\notin P_{j}$ by Lemma 4.3. We now use the fact that if $M$ is some perfect matching of the Majorana operators for a system, then $\frac{1}{\text{tr}(I)}\prod_{(k,l)\in M}(I\pm i\gamma_{k}\gamma_{l})$ is a Gaussian state [25]. Thus, each new fermionic mode added to $\sigma_{j}$ adds pairs of disjoint Majorana operators, and therefore each $\sigma_{j}$ will be an unnormalized Gaussian state, with a scaling factor obtained by various iterations of step $(g)$. ∎
The above corollary importantly does not imply that the normalization factors of these Gaussian states are nonnegative.
Corollary 4.5
The operator $\Gamma_{2n}=0$, so that $(I+\alpha_{2n}\Gamma_{2n})\sigma_{2n}=\sigma_{2n}$.
Proof.
Note that $|T_{2n}|=|T_{2n}-S_{2n}|\leq 1$ by Lemma 4.3, and the support of $\Gamma_{2n}$ is even (it is contained in $\mathcal{M}^{*}$), so $T_{2n}=\emptyset$. It follows that $\Gamma_{2n}\in\{\pm I,0\}$, and therefore $\Gamma_{2n}=0$ by step (g). ∎
Corollary 4.6
Each operator $(I+\alpha_{j}\Gamma_{j})\sigma_{j}$ is Hermitian.
Proof.
Since $T_{j}\cap P_{j}=\emptyset$, $\sigma_{j}$ and $I+\alpha_{j}\Gamma_{j}$ commute. By Corollary 4.4, $\sigma_{j}$ is Hermitian, $\alpha_{j}$ is real, and each $\Gamma_{j}$ is Hermitian, so the corollary follows. ∎
Lemma 4.7
The output distribution of the algorithm has expectation $e^{-\beta H}$.
Proof.
In order to prove this result, we will show that
$\mathbb{E}[(I+\alpha_{j+1}\Gamma_{j+1})\sigma_{j+1}|\mathcal{F}_{j}]=M_{j+1}(I+\alpha_{j}\Gamma_{j})\sigma_{j}M_{j+1}^{\dagger}.$ (19)
This result can be rewritten as
$\mathbb{E}[M_{j+1}^{-1}(I+\alpha_{j+1}\Gamma_{j+1})\sigma_{j+1}(M_{j+1}^{\dagger})^{-1}|\mathcal{F}_{j}]=(I+\alpha_{j}\Gamma_{j})\sigma_{j},$ (20)
since $M_{j+1}$ is defined deterministically in terms of $\Gamma_{j}$ and $S_{j}$, and so $M_{j+1}$ is $\mathcal{F}_{j}$-measurable. The lemma will follow from this equation, as follows. We may write:
$\mathbb{E}\left[M_{1}^{-1}\dots M_{2n}^{-1}(I+\alpha_{2n}\Gamma_{2n})\sigma_{2n}(M_{2n}^{{\dagger}})^{-1}\dots(M_{1}^{{\dagger}})^{-1}\right]$
$=\mathbb{E}\left[\mathbb{E}\left[M_{1}^{-1}\dots M_{2n-1}^{-1}M_{2n}^{-1}(I+\alpha_{2n}\Gamma_{2n})\sigma_{2n}(M_{2n}^{{\dagger}})^{-1}(M_{2n-1}^{\dagger})^{-1}\dots(M_{1}^{{\dagger}})^{-1}|\mathcal{F}_{2n-1}\right]\right]$
$=\mathbb{E}\left[M_{1}^{-1}\dots M_{2n-1}^{-1}\mathbb{E}\left[M_{2n}^{-1}(I+\alpha_{2n}\Gamma_{2n})\sigma_{2n}(M_{2n}^{{\dagger}})^{-1}|\mathcal{F}_{2n-1}\right](M_{2n-1}^{\dagger})^{-1}\dots(M_{1}^{{\dagger}})^{-1}\right]$
$=\mathbb{E}\left[M_{1}^{-1}\dots(I+\alpha_{2n-1}\Gamma_{2n-1})\sigma_{2n-1}\dots(M_{1}^{{\dagger}})^{-1}\right]$
$=\dots=\mathbb{E}\left[I+\alpha_{0}\Gamma_{0}\right]$
$=I.$
Here, we use the law of total expectation in the first equality, use that $M_{1},\dots,M_{2n-1}$ are $\mathcal{F}_{2n-1}$-measurable in the second equality, and apply equation 20 in the third equality. Repeating this $2n$ times for each $\mathcal{F}_{j}$ until $\mathcal{F}_{0}$ yields the final expression. Since $M_{1}^{-1}\dots M_{2n}^{-1}=e^{\beta H/2}$, the above formula implies that
$\mathbb{E}[e^{\beta H/2}(I+\alpha_{2n}\Gamma_{2n})\sigma_{2n}e^{\beta H/2}]=I,$
and since $e^{\beta H/2}$ is a constant we conclude that $\sigma_{2n}=(I+\alpha_{2n}\Gamma_{2n})\sigma_{2n}=e^{-\beta H}$.
To obtain equation 19, we first calculate the conditional expectation of $\left(I+\Gamma_{j+1}^{(d)}\right)\sigma_{j+1}^{(d)}$ with respect to $\mathcal{F}_{j+1}^{(c)}$:
$\mathbb{E}\left[\left(I+\Gamma_{j+1}^{(d)}\right)\sigma_{j+1}^{(d)}|\mathcal{F}_{j+1}^{(c)}\right]$ $=\Bigg{(}\left(1-\frac{1}{\mathcal{R}}\right)\left(I+\left(1-\frac{1}{\mathcal{R}}\right)^{-1}\alpha_{j}\Gamma_{j}\right)$ (21)
$\hskip 14.22636pt+\frac{1}{6\mathcal{R}}(I+6\mathcal{R}\beta_{1}\Lambda_{1})$
$\hskip 14.22636pt+\frac{1}{6\mathcal{R}}(I+6\mathcal{R}\beta_{2}\Lambda_{2})$
$\hskip 14.22636pt+\frac{1}{6\mathcal{R}}(I+6\mathcal{R}\beta_{1}\alpha_{j}\Lambda_{1}\Gamma_{j})$
$\hskip 14.22636pt+\frac{1}{6\mathcal{R}}(I+6\mathcal{R}\alpha_{j}\beta_{2}\Gamma_{j}\Lambda_{2})$
$\hskip 14.22636pt+\frac{1}{6\mathcal{R}}(I+6\mathcal{R}\beta_{1}\beta_{2}\Lambda_{1}\Lambda_{2})$
$\hskip 14.22636pt+\frac{1}{6\mathcal{R}}(I+6\mathcal{R}\beta_{1}\alpha_{j}\beta_{2}\Lambda_{1}\Gamma\Lambda_{2})\Bigg{)}\sigma_{j}.$
Grouping together the $I$ terms and simplifying, we obtain the inner term as
$I+\alpha_{j}\Gamma_{j}+\beta_{1}\Lambda_{1}+\beta_{2}\Lambda_{2}+(\beta_{1}\Lambda_{1})(\alpha_{j}\Gamma_{j})+(\alpha_{j}\Gamma_{j})(\beta_{2}\Lambda_{2})+(\beta_{1}\Lambda_{1})(\beta_{2}\Lambda_{2})+(\beta_{1}\Lambda_{1})(\alpha_{j}\Gamma_{j})(\beta_{2}\Lambda_{2}).$
Including $\sigma_{j}$ in the expression yields
$(I+\beta_{1}\Lambda_{1})(I+\alpha_{j}\Gamma_{j})(I+\beta_{2}\Lambda_{2})\sigma_{j}.$
Using the law of total expectation, we may now take the expectation of $\left(I+\alpha^{(d)}_{j+1}\Gamma_{j+1}^{(d)}\right)\sigma_{j+1}^{(d)}$ with respect to $\mathcal{F}_{j}$, obtaining
$\mathbb{E}\left[\left(I+\alpha^{(d)}_{j+1}\Gamma_{j+1}^{(d)}\right)\sigma_{j+1}^{(d)}|\mathcal{F}_{j}\right]$ $=\mathbb{E}[(I+\beta_{1}\Lambda_{1})](I+\alpha_{j}\Gamma_{j})\mathbb{E}[(I+\beta_{2}\Lambda_{2})]\sigma_{j}$ (22)
$=M_{j+1}(I+\alpha_{j}\Gamma_{j})M_{j+1}^{\dagger}\sigma_{j}$
$=M_{j+1}(I+\alpha_{j}\Gamma_{j})\sigma_{j}M_{j+1}^{\dagger}$
where the last equality holds since $S_{j}\cap P_{j}=\emptyset$, so $\sigma_{j}$ commutes with $M_{j+1}$.
In step (e), every non-Hermitian Majorana operator is set to 0. However, the operators that are not Hermitian are anti-Hermitian. Using this fact, we can expand
$\left(I+\alpha^{(e)}_{j+1}\Gamma_{j+1}^{(e)}\right)\sigma_{j+1}^{(e)}$ $=\left(I+\frac{1}{2}\alpha_{j+1}^{(d)}\left(\Gamma_{j+1}^{(d)}+\left(\Gamma_{j+1}^{(d)}\right)^{\dagger}\right)\right)\sigma_{j+1}^{(d)}$ (23)
$=\frac{1}{2}\left(\left(I+\alpha^{(d)}_{j+1}\Gamma_{j+1}^{(d)}\right)\sigma_{j+1}^{(d)}+\left(\left(I+\alpha^{(d)}_{j+1}\Gamma_{j+1}^{(d)}\right)\sigma_{j+1}^{(d)}\right)^{\dagger}\right)$
The first equality holds because $\frac{1}{2}(\Gamma+\Gamma^{\dagger})=\Gamma$ if $\Gamma$ is Hermitian and 0 if $\Gamma$ is anti-Hermitian. The second equality is satisfied because $\sigma_{j+1}^{(d)}$ is Hermitian, and commutes with $\Gamma_{j+1}^{(d)}$ since $\text{supp}\left(\Gamma_{j+1}^{(d)}\right)\subseteq S_{j}\cup T_{j}$, so it has disjoint support from $\sigma_{j+1}^{(d)}=\sigma_{j}$ by Lemma 4.3. Now, taking the conditional expectation with respect to $\mathcal{F}_{j}$, we find
$\mathbb{E}\left[\left(I+\alpha^{(e)}_{j+1}\Gamma_{j}^{(e)}\right)\sigma_{j+1}^{(e)}|\mathcal{F}_{j}\right]$ (24)
$=\frac{1}{2}\mathbb{E}\left[\left(I+\alpha^{(d)}_{j+1}\Gamma_{j+1}^{(d)}\right)\sigma_{j+1}^{(d)}|\mathcal{F}_{j}\right]+\frac{1}{2}\mathbb{E}\left[\left(I+\alpha^{(d)}_{j+1}\Gamma_{j+1}^{(d)}\right)\sigma_{j+1}^{(d)}|\mathcal{F}_{j}\right]^{\dagger}$
$=\frac{1}{2}M_{j+1}(I+\alpha_{j}\Gamma_{j})\sigma_{j}M_{j+1}^{\dagger}+\frac{1}{2}\left(M_{j+1}(I+\alpha_{j}\Gamma_{j})\sigma_{j}M_{j+1}^{\dagger}\right)^{\dagger}$
$=M_{j+1}(I+\alpha_{j}\Gamma_{j})\sigma_{j}M_{j+1}^{\dagger},$
where the last equality holds since $(I+\alpha_{j}\Gamma_{j})\sigma_{j}$ is Hermitian by Corollary 4.6.
Now, consider $I+\alpha^{(f)}_{j+1}\Gamma_{j+1}^{(f)}$. Here, we may write
$\mathbb{E}\left[\left(I+\alpha^{(f)}_{j+1}\Gamma_{j+1}^{(f)}\right)\sigma_{j+1}^{(f)}|\mathcal{F}_{j+1}^{(e)}\right]$ (25)
$=\frac{1}{2}\left(I+\alpha^{(e)}_{j+1}\Gamma^{-}_{j+1}\right)(I+i\gamma_{k}\gamma_{l})\sigma_{j+1}^{(e)}+\frac{1}{2}\left(I-\alpha_{j+1}^{(e)}\Gamma^{-}_{j+1}\right)(I-i\gamma_{k}\gamma_{l})\sigma_{j+1}^{(e)}$
$=(I+\alpha_{j+1}^{(e)}\Gamma^{-}_{j+1}(i\gamma_{k}\gamma_{l}))\sigma_{j+1}^{(e)}$
$=\left(I+\alpha_{j+1}^{(e)}\Gamma_{j+1}^{(e)}\right)\sigma_{j+1}^{(e)}.$
Finally,
$\left(I+\alpha_{j+1}\Gamma_{j+1}\right)\sigma_{j+1}$ $=\left(I+0\right)\left(I+\alpha^{(f)}_{j+1}\Gamma_{j+1}^{(f)}\right)\sigma_{j+1}^{(f)}$ (26)
$=\left(I+\alpha^{(f)}_{j+1}\Gamma_{j+1}^{(f)}\right)\sigma_{j+1}^{(f)}$
Chaining these equalities and using the law of total expectation, we obtain the desired formula,
$\mathbb{E}\left[\left(I+\alpha_{j+1}\Gamma_{j+1}\right)\sigma_{j+1}|\mathcal{F}_{j}\right]$ $=\mathbb{E}\left[\left(I+\alpha_{j+1}^{(e)}\Gamma_{j+1}^{(e)}\right)\sigma_{j+1}^{(e)}|\mathcal{F}_{j}\right]$ (27)
$=M_{j+1}(I+\alpha_{j}\Gamma_{j})\sigma_{j}M_{j+1}^{\dagger}.$
Lemma 4.8
Say that the parameter $N$ of the structural algorithm satisfies $N\geq 2$. Then, for any $j$, $\alpha_{j}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j}|-b_{j}}$, where $b_{j}=|T_{j}-S_{j}|$.
Proof.
In the base case $\alpha_{0}=0$, which clearly satisfies the hypothesis. For the inductive step, assume that $\alpha_{j}\leq(1-\frac{1}{\mathcal{R}})^{|T_{j}|-b_{j}}$. We may assume that step $(e)$ and step $(g)$ are not taken, since they set $\alpha_{j+1}$ to zero, in which case the inductive step is immediate.
The value of $\alpha_{j+1}$ is therefore decided in step (d). If the update (i) in (d) is taken, then
$\alpha_{j+1}=\left(1-\frac{1}{\mathcal{R}}\right)^{-1}\alpha_{j}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j}|-b_{j}-1}.$ (28)
If $b_{j}=1$ at the beginning of the step, then $|T_{j+1}|=|\Gamma^{-}_{j+1}|\leq\left|\Gamma^{(d)}_{j+1}\right|-2=|\Gamma_{j}|-2$ since $\Gamma^{(d)}_{j+1}$ has two Majorana operators with indices $i,k\in S_{j+1}^{c}$ that are factored out together. These operators are factored out precisely because $\Gamma_{j+1}^{(d)}=\Gamma_{j}$ when update $(i)$ is taken, so they have the same support. Since $b_{j}=1$ there was already one index $k\in T_{j}-S_{j}$. Moreover, $T_{j}$ always has even cardinality, so there must exist some index (taken to be minimal) $i\in T_{j}\cap S_{j}^{c}$. Since $S_{j+1}=S_{j}-\{i\}$ and $i$ is in the support of $\Gamma_{j+1}^{(d)}=\Gamma_{j}$, we have $i,k\in\text{supp}\left(\Gamma_{j+1}^{(d)}\right)-S_{j+1}$ and so both are factored out in step $(f)$. Now $b_{j+1}=0$, since the only two elements in $\text{supp}\left(\Gamma_{j+1}^{(d)}\right)-S_{j+1}$ were removed from the support of $\Gamma_{j}$ to produce $\Gamma_{j+1}$. It follows that $|T_{j+1}|-b_{j+1}\leq|T_{j}|-b_{j}-1$, so
$\alpha_{j+1}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j+1}|-b_{j+1}}$
as desired. If $\Gamma$ already has empty support, then $\alpha$ is set to 0 in step (g). On the other hand, if $b_{j}=0$, then since $i$ is removed from $S_{j}$ in step (a) and $\Gamma_{j+1}=\Gamma_{j}$, $b_{j+1}=1$. It follows that $\alpha_{j+1}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j+1}|-b_{j+1}}$ as desired.
We now consider the remaining possible update steps. For (ii), (iii), note that $\frac{6\mathcal{R}}{N\mathcal{R}}=\frac{24}{N}\cdot\frac{1}{4}\leq\frac{24}{N}(1-\frac{1}{\mathcal{R}})^{\mathcal{R}}$ for $\mathcal{R}\geq 2$, as we have assumed. Moreover, since $N\geq 2$, $(N\mathcal{R})^{-k}\leq 4^{-k}\leq(1-\frac{1}{\mathcal{R}})^{\mathcal{R}k}$. Together, these statements demonstrate that $(6\mathcal{R})(N\mathcal{R})^{-k}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{\mathcal{R}k}$, and so in cases (ii) and (iii) we may deduce that
$\alpha_{j+1}=6\mathcal{R}(N\mathcal{R})^{-t_{1}}$ $\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{\mathcal{R}t_{1}}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|\Lambda_{1}|},$ (29)
$\alpha_{j+1}=6\mathcal{R}(N\mathcal{R})^{-t_{2}}$ $\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{\mathcal{R}t_{2}}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|\Lambda_{2}|}.$
since $|\Lambda_{1}|\leq\mathcal{R}t_{1}$ and $|\Lambda_{2}|\leq\mathcal{R}t_{2}$ by Lemma 3.2. Since in cases (ii) and (iii), $\Gamma_{j+1}$ has support contained in $\text{supp}(\Lambda_{1})$ and $\text{supp}(\Lambda_{2})$ respectively, we conclude
$\alpha_{j+1}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j+1}|}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j+1}|-b_{j+1}}.$ (30)
Similarly, for case (vi), using the inequality
$\alpha_{j+1}$ $=6\mathcal{R}(N\mathcal{R})^{-(t_{1}+t_{2})}$ (31)
$\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{\mathcal{R}(t_{1}+t_{2})}$
$\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|\Lambda_{1}\Lambda_{2}|}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j+1}|-b_{j+1}}$
Cases (iv), (v), and (vii) remain. These three cases are dealt with very similarly, so we only reason through case (vii).
There are two possibilities worth considering. In the first, $b_{j+1}\geq b_{j}$. Here, we write that
$\alpha_{j+1}$ $\leq 6\mathcal{R}(N\mathcal{R})^{-(t_{1}+t_{2})}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j}|-b_{j}}$ (32)
$\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j}|-b_{j}+\mathcal{R}(t_{1}+t_{2})}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j+1}|-b_{j+1}},$
since $|T_{j+1}|=|\Gamma_{j+1}|\leq|\Lambda_{1}\Gamma_{j}\Lambda_{2}|\leq|T_{j}|+\mathcal{R}(t_{1}+t_{2})$.
Now, say that $b_{j+1}<b_{j}$, i.e. $b_{j+1}=0$ and $b_{j}=1$. If we say $\{i\}=S_{j}-S_{j+1}$, then we know that $i\in S_{j+1}^{c}$, and therefore $i\notin\Gamma_{j+1}$, or else it would be true that $i\in T_{j+1}-S_{j+1}$ and $b_{j+1}>0$. We may therefore remark that
$|\Gamma_{j+1}|\leq|\Lambda_{1}|+|\Gamma_{j}|+|\Lambda_{2}|-1\leq|T_{j}|+\mathcal{R}(t_{1}+t_{2})-1.$ (33)
The first inequality arises precisely because $\Gamma_{j+1}$ has support contained in the union of $\Lambda_{1}$, $\Gamma_{j}$, and $\Lambda_{2}$, but also does not contain the index $i\in T_{j}$.
Just as for the first possibility, it now follows that
$\alpha_{j+1}$ $\leq 6\mathcal{R}(N\mathcal{R})^{-(t_{1}+t_{2})}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j}|-b_{j}}$
$\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j}|-b_{j}-1+\mathcal{R}(t_{1}+t_{2})}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j+1}|-b_{j+1}}.$
We may conclude the inductive hypothesis no matter which choice is taken in step (d), completing the proof. ∎
Lemma 4.9
For any $\beta\leq\frac{1}{48\mathcal{R}d}$, and choosing $N=24$ in the structural algorithm, the trace of any output state $\sigma$ of the algorithm satisfies $\text{tr}(\sigma)\geq 0$.
Proof.
Since $\beta\leq\frac{1}{48\mathcal{R}d}=\frac{1}{2N\mathcal{R}d}$, the previous lemma applies. To prove this result, it suffices to prove that whenever step $(g)$ of the algorithm is undertaken, $\alpha_{j}\leq 1$, so that the scaling factor $I+\alpha_{j}\Gamma_{j}$ is nonnegative. Indeed, other than these factors, $\sigma_{2n}$ can be expressed as $\Pi_{(k,l)\in M}(I\pm i\gamma_{k}\gamma_{l})$. Expanding out this product, since the pairs $(k,l)$ do not overlap, every summand has nonzero support except $I$, and so
$\text{tr}\left(\Pi_{(k,l)\in M}(I\pm i\gamma_{k}\gamma_{l})\right)=\text{tr}(I)\geq 0.$
If each of the step $(g)$ factors is nonnegative, we obtain the desired property $\text{tr}(\sigma_{2n})\geq 0$. Now, we can simply apply Lemma 4.8. When time step $(g)$ is taken, $|T_{j}|=|\Gamma_{j}|=0$ and therefore $b_{j}=0$ as well. So the lemma implies that
$\alpha_{j}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j}|-b_{j}}=\frac{24}{N}=1.$ (34)
Corollary 4.10
For $\beta\leq\frac{1}{48\mathcal{R}d}$, The Gibbs state $\frac{e^{-\beta H}}{\text{tr}(e^{-\beta H})}$ is a convex combination of Gaussian states.
Proof.
By Corollary 4.4, Lemma 4.7, and Lemma 4.9, $e^{-\beta H}$ is a convex combination of unnormalized Gaussian states. This implies that the Gibbs state is a convex combination of Gaussian states. ∎
5 Sampling Algorithm
5.1 Overview
In this section, we upgrade the structural algorithm in Section 4 to a sampling algorithm that efficiently samples from the Gibbs state of a $\mathcal{R}$-local, degree-$d$ fermionic Hamiltonian at sufficiently high temperature. Specifically, when the inverse temperature satisfies
$\beta\leq\frac{1}{50\mathcal{R}d^{2}},$
we show that a classical algorithm can efficiently sample from a distribution of fermionic Gaussian states, such that the expectation approximates the Gibbs state
$\rho_{\beta}=\frac{e^{-\beta H}}{\text{tr}(e^{-\beta H})}$
up to trace distance $\epsilon$. This extends the structural result in Section 4 and shows that Gibbs sampling is classically easy in the high-temperature fermionic regime as well.
We will first comment on the reduced temperature threshold needed for the sampling algorithm, detailed in Section 5.2. First, we slightly tighten the temperature bound compared to Section 4, from $\frac{1}{48\mathcal{R}^{2}}$ to $\frac{1}{50\mathcal{R}^{2}}$, to ensure that the coefficients $\alpha$’s arising in the telescoping expansion is bounded away from $1$. Previously, only $\alpha\leq 1$ was needed, to ensure that the intermediate operators in the form of $I+\alpha\Gamma$ remained positive semidefinite. In Lemma 5.1 below, this bound is amplified slightly. Additionally, we now need to efficiently approximate the partition function $\text{tr}(e^{-\beta H})$, which is necessary to normalize samples. This introduces another factor of $d^{-1}$ needed for the temperature threshold, as shown in Corollary 5.3.
In order to design an efficient sampling algorithm, we reinterpret the structural algorithm from Section 4 as implicitly generating a sampling tree $\mathcal{T}$, introduced in Section 5.3. Each node in $\mathcal{T}$ represents a step of the sampling process, and each edge corresponds to a stochastic outcome of that step, arising from the random choices of sampling an operator from the expansion of $M$ or $M^{\dagger}$, or performing an update or pinning step. A complete path from the root to a leaf in $\mathcal{T}$ specifies a sequence of random choices leading to an unnormalized fermionic Gaussian state, and each path carries a weight determined by the product of probabilities along the path. The distribution over all leaves encodes a convex decomposition of $e^{-\beta H}.$ However, this is a distribution of unnormalized Gaussian states with varying normalization constants. The algorithm that samples from $\mathcal{T}$ with a distribution that must be corrected by the normalization constants. $\mathcal{T}$ is also not efficient to sample from, for instance because nodes can have infinitely many children. We show that by simulating a well-chosen random walk (Theorem 5.5) on an approximate version of $\mathcal{T}$, $\mathcal{T}^{\prime}$, an efficient sampling algorithm for the corrected distribution may be obtained. $\mathcal{T}^{\prime}$ is defined in Section 5.4. Analyzing the performance of this random walk, we show that high-temperature fermionic Gibbs states can be efficiently sampled.
5.2 Temperature Threshold
In the sampling setting, the bound on the inverse temperature $\beta$ changes from the structural algorithm by a factor of $\frac{24}{25}d^{-1}$. The reason for the factor of $\frac{24}{25}$ is the following variant of Lemma 4.8:
Lemma 5.1
For $\beta\leq\frac{1}{50\mathcal{R}d^{2}}$ and for the structural algorithm with $N=25$, $\alpha_{j}\leq\frac{24}{25}$, for any $j$.
Proof.
Since $\beta\leq\frac{1}{50\mathcal{R}d^{2}}\leq\frac{1}{50\mathcal{R}d}=\frac{1}{2N\mathcal{R}d}$, Lemma 4.8 implies that
$\alpha_{j}\leq\frac{24}{N}\left(1-\frac{1}{\mathcal{R}}\right)^{|T_{j}|-b_{j}}$
Note that since $b_{j}=|T_{j}-S_{j}|$, $|T_{j}|-b_{j}\geq 0$. It follows that $\alpha_{j}\leq\frac{24}{25}$. ∎
To use the structural algorithm as a subroutine for sampling, it will be necessary that $\alpha_{j}$ is bounded away from 1, which is why the choice of a slightly larger temperature bound ensures that $\alpha_{j}\leq 1-\frac{1}{25}$.
The additional factor of $d^{-1}$ allows for the efficient approximation of the partition function of the Hamiltonians $H^{S}$ for any $S\subseteq[n]$. To do so, we cite Theorem 3.7 from [5]. We remark that this theorem, although originally proved for the spin setting, can be applied without modification to the fermionic setting because the terms in the fermionic Hamiltonian, having even parity, are still “bosonic” in the sense that non-overlapping terms commute with each other.
Lemma 5.2
Say $H$ is an $n$-mode fermionic Hamiltonian with fixed locality $\mathcal{R}$ and degree $d$. For $\beta<\frac{1}{e(e+1)(1+e(d-1))d}$, the partition function $\text{tr}(e^{-\beta H})$ may be estimated to multiplicative precision $\epsilon$ in time $\text{poly}(n,\epsilon^{-1})$.
Corollary 5.3
When $\beta\leq\frac{1}{50\mathcal{R}d^{2}}$, the partition function of an $n$-mode fermionic Hamiltonian $H$ with fixed locality $\mathcal{R}$ and degree $d$ as well as the quantity $\text{tr}(e^{-\beta H})$ can be estimated to multiplicative precision $\epsilon$ in time $\text{poly}(n,\epsilon^{-1})$.
Proof.
For $\mathcal{R}\geq 2$,
$\frac{1}{50\mathcal{R}d^{2}}\leq\frac{1}{100d^{2}}<\frac{1}{e(e+1)(1+e(d-1))d}$ (35)
for any $d\geq 1$. ∎
For this choice of $\beta$, therefore, we may use the partition function and the structural algorithm together to implement a Gibbs sampling algorithm. In particular, we will make use of a subroutine of the algorithm, which is essentially just one of the $2n$ steps in Section 4.2 with $N=25$:
Algorithm 5.4 (Pinning one site)
Input: Subset of terms $S\subseteq[2n]$ for unpinned sites and iteration $j$. Pinned unnormalized Gaussian states $\sigma$ and unpinned state $I+\alpha\Gamma.$
Output: $(j^{\prime},S^{\prime},\sigma^{\prime},\Gamma^{\prime},\alpha^{\prime})$, representing the pinned version of the input states at the next iteration.
  1. 1.
    If $\text{supp}(\Gamma)\cap S$ is nonempty, set $i$ as the minimal index in $\text{supp}(\Gamma)\cap S$. Otherwise, let $i$ be the minimal index in $S$. Define $S^{\prime}=S-\{i\}$. Define
    $M=e^{\beta H^{S^{\prime}/2}}e^{-\beta H^{S/2}}.$
  2. 2.
    Sample $I+\beta_{1}\Lambda_{1}$ for $\beta_{1}=(25\mathcal{R})^{-t_{1}}$ from the probability distribution in Lemma 3.2, so that
    $\mathbb{E}\left[I+\beta_{1}\Lambda_{1}\right]=M=e^{\beta H^{S^{\prime}/2}}e^{-\beta H^{S/2}}.$
  3. 3.
    Sample $I+\beta_{2}\Lambda_{2}$ for $\beta_{2}=(25\mathcal{R})^{-t_{2}}$ independently, so that
    $\mathbb{E}\left[I+\beta_{2}\Lambda_{2}\right]=M^{\dagger}.$
  4. 4.
    Set $\sigma^{\prime}\leftarrow\sigma$ and $(\Gamma^{\prime},\alpha^{\prime})$ as:
    1. (a)
      $\bigl{(}\Gamma,(1-\frac{1}{\mathcal{R}})^{-1}\alpha_{j}\bigl{)}$ with probability $1-\frac{1}{\mathcal{R}}$
    2. (b)
      $\left(\Lambda_{1},(6\mathcal{R})\beta_{1}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    3. (c)
      $\left(\Lambda_{2},(6\mathcal{R})\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    4. (d)
      $\left(\Lambda_{1}\Gamma,(6\mathcal{R})\beta_{1}\alpha_{j}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    5. (e)
      $\left(\Gamma\Lambda_{2},(6\mathcal{R})\alpha_{j}\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    6. (f)
      $\left(\Lambda_{1}\Lambda_{2},(6\mathcal{R})\beta_{1}\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    7. (g)
      $\left(\Lambda_{1}\Gamma\Lambda_{2},(6\mathcal{R})\beta_{1}\alpha_{j}\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
  5. 5.
    If $\Gamma^{\prime}$ is not Hermitian:
    • $\Gamma^{\prime}\leftarrow 0,\alpha^{\prime}\leftarrow 0$.
  6. 6.
    If there are two indices $k,l\in S_{j}^{c}$ such that $k,l\in\text{supp}\left(\Gamma^{\prime}\right)$, i.e. $\Gamma^{\prime}=\Gamma^{-}\cdot(i\gamma_{k}\gamma_{l})$:
    1. (a)
      $\Gamma^{\prime}\leftarrow\Gamma^{-}$, $\sigma^{\prime}\leftarrow(I+i\gamma_{k}\gamma_{l})\cdot\sigma^{\prime}$ with probability $\frac{1}{2}$
    2. (b)
      $\Gamma^{\prime}\leftarrow-\Gamma^{-}$, $\sigma^{\prime}\leftarrow(I-i\gamma_{k}\gamma_{l})\cdot\sigma^{\prime}$ with probability $\frac{1}{2}$
  7. 7.
    If $\Gamma^{\prime}\in\{\pm I\}$:
    • $\sigma^{\prime}\leftarrow(I+\alpha^{\prime}\Gamma^{\prime})\sigma^{\prime}$, $\Gamma^{\prime}\leftarrow 0$, and $\alpha^{\prime}\leftarrow 0$.
  8. 8.
    $j\leftarrow j+1$.
Recall that the structural algorithm samples unnormalized Gaussian states $\sigma_{i}$ with probability $p_{i}$ such that
$\sum_{i}p_{i}\sigma_{i}=e^{-\beta H}.$
As discussed in the overview, a classical sampling algorithm must sample the operators $\rho_{i}=\frac{\sigma_{i}}{\text{tr}(\sigma_{i})}$ according to the distribution
$q_{i}=\frac{p_{i}\text{tr}(\sigma_{i})}{\text{tr}(e^{-\beta H})}.$
Preparing the operators $\rho_{i}$ with respect to the distribution $p_{i}$ is easy, because the classical description of $\sigma_{i}$ can easily be renormalized to obtain a classical description of $\rho_{i}$. Thus, the main obstacle is converting this naive sampler to one that can sample $\rho_{i}$ with the correct probabilities, $q_{i}$, yielding a classical algorithm that can prepare
$\sum_{i}q_{i}\rho_{i}=\frac{e^{-\beta H}}{\text{tr}(e^{-\beta H})}.$
5.3 Sample Tree
In general, reducing a sampler for a distribution $q$ to a sampler for a distribution $p$ is a daunting task. However, there is a natural tree structure, $\mathcal{T}$, associated with these distributions. The tree has a root node with associated data $(0,S_{0},\sigma_{0},\Gamma_{0},\alpha_{0})=(0,[2n],I,0,0)$, and the children of a given node correspond to all possible random choices in the Algorithm 5.4, with the algorithm’s output $(j+1,S_{j+1},\sigma_{j+1},\Gamma_{j+1},\alpha_{j+1})$ as the associated data. For a node $v$, we will denote its associated quantities by $j_{v},S_{v},\sigma_{v},\Gamma_{v}$, and $\alpha_{v}$. For any internal node $v$ and any (potentially unnormalized) distribution on the leaves $r$, we use the notation $\mathcal{L}_{\mathcal{T}}(v)$ to denote the leaves below a given node $v$, and $r_{v}=\sum_{w\in\mathcal{L}_{T}(v)}r_{w}$. Our structural algorithm is a walk down this tree from $j=0$ to $j=2n$, by repeated application of Algorithm 5.4, and the probability distribution $p$ on the leaves is the probability of reaching a given leaf applying this algorithm. $p_{v}$ for a general internal node is therefore the probability of reaching the node $v$. One final piece of notation is that $\mathcal{C}_{\mathcal{T}}(v)$ will be used to denote the set of children of a given node $v$.
A reduction between sampling algorithms of two distributions on the leaves of a tree is possible when additional conditions are satisfied on the internal nodes of the tree. In particular, to convert a sampling algorithm for $p_{i}$ to a sampling algorithm for $q_{i}$, we make use of Theorem 5.8 in [5]. We replace a certain approximation ratio in their formulation, 10, with an arbitrary constant $\gamma$. We note that proof of the theorem in [5] generalizes to this case as well.
Theorem 5.5
Assume that $p$ and $q$ are two probability distributions on the leaves of a tree $\mathcal{T}$ of depth $d$, where each node has at most $k$ children. Let $\hat{p}$ and $\hat{q}$ be any unnormalized distributions that are equal to $p$ and $q$ up to some constant, i.e. $\hat{p}_{v}=Cp_{v}$ and $\hat{q}_{v}=Dq_{v}$ for any node $v$ and some constants $C,D\neq 0$. Assume also the following four conditions:
  1. 1.
    For any internal node $v$, the values $\frac{\hat{q}_{v}}{\hat{p}_{v}}$ can be calculated efficiently up to constant multiplicative error.
  2. 2.
    For any leaf $v$, the values $\frac{\hat{q}_{v}}{\hat{p}_{v}}$ can be efficiently calculated exactly.
  3. 3.
    For any internal node $v$, the leaves $v_{k}$ of $v$ can be sampled efficiently according to the distribution $\frac{p_{v_{k}}}{p_{v}}$.
  4. 4.
    Finally, for any node $v$ and its child $w$,
    $\left|\log\left(\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)-\log\left(\frac{\hat{q}_{w}}{\hat{p}_{w}}\right)\right|\leq\log(\gamma)$
    for some constant $\gamma>1$.
Then, there exists a $\text{poly}\left(d,\log(k),\log\left(\epsilon^{-1}\right)\right)$ algorithm to sample from $q$ up to total variation distance $\epsilon$.
Theorem 5.5 works by performing a carefully chosen random walk on the tree $\mathcal{T}$. This walk is designed so that its marginal distribution on the leaves of the tree converges to the target distribution $q$, using only local information about the ratios $\hat{q}_{v}/\hat{p}_{v}$. At each node of the tree, the walk chooses to move up, down, or stay in place based on local comparisons of these unnormalized probabilities. The assumption that the log-ratio $\log(\hat{q}_{v}/\hat{p}_{v})$ varies slowly across parent-child pairs is crucial: if the changes were large, they could prevent the walk from effectively concentrating around the parts of the tree where $q$ is most dense. We refer readers to Algorithm 5.9 in [5] for the technical details of the walk as well as its mixing time analysis.
We will first establish that almost all of the conditions of Theorem 5.5 can be satisfied for the sampling tree $\mathcal{T}$ of Algorithm 5.4. The analysis requires an estimate of how the trace of a positive semidefinite matrix is affected by perturbations.
Lemma 5.6
Assume that $P$ is a positive semidefinite operator, and $A$ is some operator such that $\lVert A\rVert_{\infty}\leq 1-\epsilon$ for $0<\epsilon\leq 1$. Then
$\left|\log(\text{tr}(P(1+A)))-\log(\text{tr}(P))\right|\leq\text{log}\left(\epsilon^{-1}\right)$
Proof.
We may use Hölder’s inequality to say that
$|\text{tr}(PA)|$ $\leq\lVert PA\rVert_{1}$ (36)
$\leq\lVert P\rVert_{1}\lVert A\rVert_{\infty}$
$=\text{tr}(P)\lVert A\rVert_{\infty}$
using the property that the $1$-norm is the sum of the absolute values of the eigenvalues of an operator. From this, we deduce that
$\text{tr}(P)(I-\lVert A\rVert_{\infty})\leq\text{tr}(P(1+A))\leq\text{tr}(P)(I+\lVert A\rVert_{\infty}),$ (37)
and in particular by the bound on $\lVert A\rVert_{\infty}$:
$\epsilon\text{tr}(P)\leq\text{tr}(P(1+A))\leq(2-\epsilon)\text{tr}(P).$ (38)
It follows that
$\left|\log(\text{tr}(P(1+A)))-\log(\text{tr}(P))\right|\leq\max(\log(\epsilon^{-1}),\log(2-\epsilon))=\log(\epsilon^{-1}),$ (39)
where the last equality holds since $(2-\epsilon)\epsilon\leq 1$ for $0<\epsilon\leq 1$. ∎
To prove condition (1), it will also be important to track the expectation of the output of Algorithm 4.1 conditioned on reaching a given internal node of $\mathcal{T}$.
Lemma 5.7
Let $A_{v}$ be the event of reaching an internal node $v$ in the structural algorithm. Then $\mathbb{E}[\sigma_{2n}|A_{v}]=e^{-\beta H^{S_{v}}/2}(I+\alpha_{v}\Gamma_{v})\sigma_{v}e^{-\beta H^{S_{v}}/2}$.
Proof.
Say that the node $v$ corresponds to the $j$th step of the algorithm. As in Lemma 4.7, we can use Equation 20, along with the law of total expectations and properties of $\mathcal{F}$-measurable random variables:
$\mathbb{E}\left[M_{j}^{-1}\dots M_{2n}^{-1}(I+\alpha_{2n}\Gamma_{2n})\sigma_{2n}(M_{2n}^{{\dagger}})^{-1}\dots(M_{j}^{{\dagger}})^{-1}|\mathcal{F}_{j}\right]$
$=\mathbb{E}\left[\mathbb{E}\left[M_{j}^{-1}\dots M_{2n-1}^{-1}M_{2n}^{-1}(I+\alpha_{2n}\Gamma_{2n})\sigma_{2n}(M_{2n}^{{\dagger}})^{-1}(M_{2n-1}^{\dagger})^{-1}\dots(M_{j}^{{\dagger}})^{-1}|\mathcal{F}_{2n-1}\right]|\mathcal{F}_{j}\right]$
$=\mathbb{E}\left[M_{j}^{-1}\dots M_{2n-1}^{-1}\mathbb{E}\left[M_{2n}^{-1}(I+\alpha_{2n}\Gamma_{2n})\sigma_{2n}(M_{2n}^{{\dagger}})^{-1}|\mathcal{F}_{2n-1}\right](M_{2n-1}^{\dagger})^{-1}\dots(M_{j}^{{\dagger}})^{-1}|\mathcal{F}_{j}\right]$
$=\mathbb{E}\left[M_{j}^{-1}\dots(I+\alpha_{2n-1}\Gamma_{2n-1})\sigma_{2n-1}\dots(M_{j}^{{\dagger}})^{-1}|\mathcal{F}_{j}\right]$
$=\dots=\mathbb{E}\left[(I+\alpha_{j}\Gamma_{j})\sigma_{j}|\mathcal{F}_{j}\right]$
$=(I+\alpha_{j}\Gamma_{j})\sigma_{j}.$
Note that $M_{2n}\dots M_{j}=e^{-\beta H^{S_{j}}/2}$ and $\alpha_{2n}=0$, the left hand side of the above expression is
$\mathbb{E}\left[e^{\beta H^{S_{j}}/2}(I+\alpha_{2n}\Gamma_{2n})\sigma_{2n}e^{\beta H^{S_{j}}/2}|\mathcal{F}_{j}\right]=e^{\beta H^{S_{j}}/2}\mathbb{E}\left[\sigma_{2n}|\mathcal{F}_{j}\right]e^{\beta H^{S_{j}}/2},$ (40)
since $S_{j}$ depends only on randomness up until step $j$. We conclude that
$\mathbb{E}\left[\sigma_{2n}|\mathcal{F}_{j}\right]=e^{-\beta H^{S_{j}}/2}(I+\alpha_{j}\Gamma_{j})\sigma_{j}e^{-\beta H^{S_{j}}/2}$ (41)
Since $A_{v}$ is an event based only on random choices up to the $j$th step, it is true that
$\mathbb{E}[\cdot|A_{v}]=\mathbb{E}[\mathbb{E}[\cdot|\mathcal{F}_{j}]|A_{v}],$
from which it follows that
$\mathbb{E}[\sigma_{2n}|A_{v}]=e^{-\beta H^{S_{v}}/2}(I+\alpha_{v}\Gamma_{v})\sigma_{v}e^{-\beta H^{S_{v}}/2}.$
Lemma 5.8
Let $\mathcal{T}$ be the tree associated with Algorithm 5.4 with the associated distributions $p$ and $q$, and let $\hat{p}_{v}=p_{v}$, $\hat{q}_{v}=\text{tr}(e^{-\beta H})q_{v}=\text{tr}(\sigma_{v})p_{v}$ be the corresponding unnormalized distributions for any leaf $v$. Then conditions (1) and (2) of Theorem 5.5 are satisfied.
Proof.
See that
$\frac{\hat{q}_{v}}{\hat{p}_{v}}$ $=\frac{\sum_{w\in\mathcal{L}_{T}(v)}\hat{q}_{w}}{\sum_{w\in\mathcal{L}_{T}(v)}p_{w}}$ (42)
$=\frac{\sum_{w\in\mathcal{L}_{T}(v)}\text{tr}(\sigma_{w})p_{w}}{\sum_{w\in\mathcal{L}_{T}(v)}p_{w}}$ (43)
[…]
$=\text{tr}(e^{-\beta H^{S_{v}}/2}(I+\alpha_{v}\Gamma_{v})\sigma_{v}e^{-\beta H^{S_{v}}/2})$ (46)
$=\text{tr}(\sigma_{v}e^{-\beta H^{S_{v}}}(I+\alpha_{v}\Gamma_{v}))$ (47)
where Eq.  (46) uses Lemma 5.7. The state $\sigma$ acts on different modes than $e^{-\beta H^{S_{v}}}$, as we have established its suppport is contained in $S_{v}^{c}$. As a consequence,
$\text{tr}\left(\sigma_{v}e^{-\beta H^{S_{v}}}\right)=\text{tr}(\sigma_{v})\text{tr}\left(e^{-\beta H^{S_{v}}}\right).$ (48)
For the same reason, $\sigma_{v}e^{-\beta H^{S_{v}}}$ is positive semidefinite. Setting $P=\sigma_{v}e^{-\beta H^{S_{v}}}$ and $A=\alpha_{v}\Gamma_{v}$, note that $\lVert A\rVert_{\infty}=\alpha_{v}$. Applying Lemma 5.6,
$\left\lvert\log\left(\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)-\log(\text{tr}(P))\right\rvert$ $=\left\lvert\log\left(\text{tr}(P(I+A))\right)-\log(\text{tr}(P))\right\rvert$
$\leq\log((1-\alpha_{v})^{-1})$
$\leq\log(25),$
where the last inequality holds by Lemma 5.1. We note that $\text{tr}(P)=\text{tr}(\sigma_{v})\text{tr}\left(e^{-\beta H^{S_{v}}}\right)$, and $\text{tr}(\sigma_{v})$ is known exactly. Since $H^{S_{v}}$ has locality and degree bounded by the constant locality $\mathcal{R}$ and degree $d$ of $H$, Corollary 5.3 may therefore be applied to calculate its $\frac{\hat{q}_{v}}{\hat{p}_{v}}$ up to multiplicative error $25$. This establishes condition (1).
For condition (2), note that associated to any leaf $v$ is the state $\sigma_{v}$, whose classical representation includes the trace $\text{tr}(\sigma_{v})$. It follows that $\frac{\hat{q}_{v}}{\hat{p}_{v}}=\text{tr}(\sigma_{v})$ can be calculated exactly for any leaf $v$. Thus, condition (2) can be satisfied as well. ∎
Condition (3) does not hold true for the tree $\mathcal{T}$. One simple way to see this is that in every application of Algorithm 5.4, there are infinite possible outcomes of the random choices, due to the sampling from probability distributions with infinite support in steps 2 and 3. Furthermore, Condition (4) is satisfiable, using the following lemma. After establishing this condition, we will design a subtree $\mathcal{T}^{\prime}$ of $\mathcal{T}$ for which all four conditions may be attained, and ensure that the parameter $k$ in Theorem 5.5 is such that $\log(k)\in\text{poly}(n)$.
Lemma 5.9
For any $S_{i}$ and $S_{i-1}$ as in Lemma 3.1, if $\beta\leq\frac{1}{2Cd}$ for some value $C>1$, then $e^{-\beta H^{S_{i}}}e^{\beta H^{S_{i-1}}}=I+A$ where $\lVert A\rVert_{\infty}\leq\frac{1}{C-1}$.
Proof.
Lemma 3.1 decomposes $M_{i}$ as $I+\sum_{t>0}\sum_{j}c_{j,t}\Gamma_{j,t}$ such that $\sum_{j}c_{j,t}\leq C^{-t}$. Defining $A=\sum_{t>0}\sum_{j}c_{j,t}\Gamma_{j,t}$, we find that
$\lVert A\rVert_{\infty}\leq\sum_{t>0}\sum_{j}c_{j,t}\leq\sum_{t>0}C^{-t}=\frac{1}{C-1}.$ (49)
Lemma 5.10
Let $\mathcal{T}$ be the tree associated with Algorithm 5.4 with the associated distributions $p$ and $q$, and let $\hat{p}_{v}=p_{v}$, $\hat{q}_{v}=\text{tr}(e^{-\beta H})q_{v}=\text{tr}(\sigma_{v})p_{v}$ be the corresponding unnormalized distributions for any leaf $v$. Then condition (4) of Theorem 5.5 is satisfied, with constant $\gamma_{\mathcal{T}}=16000$.
Next, we verify condition (4) of Theorem 5.5, that
$\left|\log\left(\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)-\log\left(\frac{\hat{q}_{w}}{\hat{p}_{w}}\right)\right|\leq\log(\gamma_{\scriptscriptstyle\mathcal{T}}),$
for the constant $\gamma_{\scriptscriptstyle\mathcal{T}}=16000$. Given that $\frac{\hat{q}_{v}}{\hat{p}_{v}}=\text{tr}(e^{-\beta H^{S_{v}}}(I+\alpha_{v}\Gamma_{v})\sigma_{v})$, from the proof of Lemma 5.8, we need only show that
$\left|\log\left(\text{tr}(e^{-\beta H^{S_{v}}}(I+\alpha_{v}\Gamma_{v})\sigma_{v})\right)-\log\left(\text{tr}(e^{-\beta H^{S_{w}}}(I+\alpha_{w}\Gamma_{w})\sigma_{w})\right)\right|=O(1).$ (50)
By an application of the triangle inequality, the following three relations suffices:
$\left|\log\left(\text{tr}(\sigma_{v}e^{-\beta H^{S_{v}}}(I+\alpha_{v}\Gamma_{v}))\right)-\log\left(\text{tr}(\sigma_{v}e^{-\beta H^{S_{v}}})\right)\right|=O(1),$
$\left|\log\left(\text{tr}(\sigma_{w}e^{-\beta H^{S_{w}}}(I+\alpha_{w}\Gamma_{w}))\right)-\log\left(\text{tr}(\sigma_{w}e^{-\beta H^{S_{w}}})\right)\right|=O(1),$
$\left|\log\left(\text{tr}(\sigma_{v})\text{tr}(e^{-\beta H^{S_{v}}})\right)-\log\left(\text{tr}(\sigma_{w})\text{tr}(e^{-\beta H^{S_{w}}})\right)\right|=O(1).$
The first two have already been shown in Lemma 5.8 in the proof of condition (1), each with the bound $25=O(1)$.
For the third relation, applying Lemma 5.9 to $S_{v}$ and $S_{w}$ and using the triangle inequality, we may rewrite the inequality as
$\left|\log\left(\text{tr}(e^{-\beta H^{S_{w}}}(I+A))\right)-\log\left(\text{tr}(e^{-\beta H^{S_{w}}})\right)\right|+\left|\log\left(\text{tr}(\sigma_{v})\right)-\log\left(\text{tr}(\sigma_{w})\right)\right|=O(1)$
for some $A$. The second summand is at most $\log(25)$, since in any step $\sigma_{v}$ is multiplied by a commuting operator with trace 1 or trace $1\pm\alpha_{v}$. This latter value lies in $[\frac{1}{25},1+\frac{24}{25}]$, and therefore the logarithm is bounded in absolute value by $\log(25)$. Since $\beta\leq\frac{1}{50\mathcal{R}d^{2}}\leq\frac{1}{100d}$, we can take $C=25$, so $\lVert A\rVert_{\infty}\leq\frac{1}{49}$. Lemma 5.6 applies again since $e^{-\beta HS_{w}}$ is positive, yielding an upper bound of $\log(25\cdot 49\cdot 48^{-1})$, including the second summand.
In particular, we have shown that
$\left|\log\left(\text{tr}(e^{-\beta H^{S_{v}}}(I+\alpha_{v}\Gamma_{v}))\right)-\log\left(\text{tr}(e^{-\beta H^{S_{w}}}(I+\alpha_{w}\Gamma_{w}))\right)\right|$ (51)
$\leq\log\left(\left(\frac{1}{25}\right)^{-3}\cdot\left(\frac{48}{49}\right)^{-1}\right)$
$\leq\log(\gamma_{\scriptscriptstyle\mathcal{T}}),$
as desired.
5.4 Truncated Sample Tree
Theorem 5.11 (efficient algorithm)
Assume that $H$ has locality $\mathcal{R}$ and has degree $d$, and $\beta\leq\frac{1}{50\mathcal{R}d^{2}}$. Then there exists a classical algorithm that can sample the Gibbs state to trace distance $\epsilon$ in time $\text{poly}(n,\log(\epsilon^{-1}))$.
To prove this theorem, we will create an approximate version of Algorithm 5.4, such that the corresponding tree $\mathcal{T}^{\prime}$ will satisfy the conditions of Theorem 5.5. The following lemmas yields an approximate version of a sampling step in the algorithm.
Lemma 5.12
There exists a $\text{poly}\left(\delta^{-1}\right)$-time algorithm that samples from a distribution $\mu^{\prime}$ of outcomes $(t_{k},I+(25\mathcal{R})^{-t_{k}}\Gamma_{k})$ of Lemma 3.2, which is precisely the distribution $\mu$ from Lemma 3.2 conditioned on an event $A$ with $\mathbb{P}_{\mu}[A]\geq 1-\delta$, i.e.
$\mu^{\prime}_{X}=\begin{cases}\displaystyle\frac{\mu_{X}}{\mathbb{P}_{\mu}[A]}&\text{if }X\in A\\ 0&\text{if }X\notin A\end{cases}.$ (52)
Specifically, $A$ is the event that $t_{k}\leq t_{\text{max}}$ for threshold $t_{\text{max}}=\lceil\log_{2}\left(\delta^{-1}\right)\rceil$. Moreover, the distribution $\mu^{\prime}$ has support with size $\text{exp}(\text{polylog}\left(\delta^{-1}\right))$.
Proof.
The distribution $\mu$ corresponds to the probabilities in the convex combination
$M_{i}=\sum_{t>0}\left(\left(2^{-t}-c_{t}(N\mathcal{R})^{t}\right)I+\sum_{j}c_{j,t}(N\mathcal{R})^{t}(I+(N\mathcal{R})^{-t}\Gamma_{j,t})\right),$ (53)
where $\Gamma_{j,t}$ are $\mathcal{R}t$-local Majorana operators and $c_{j,t}\in\mathbb{R}^{+}.$ We will show that the probabilities in the convex combination
$\sum_{t=1}^{t_{\text{max}}}\left(\frac{\left(2^{-t}-c_{t}(N\mathcal{R})^{t}\right)}{1-2^{-t_{\text{max}}}}I+\sum_{j}\frac{c_{j,t}(N\mathcal{R})^{t}}{1-2^{-t_{\text{max}}}}(I+(N\mathcal{R})^{-t}\Gamma_{j,t})\right)$ (54)
can be sampled efficiently, and this is precisely the original distribution conditioned on the event $t_{k}\leq t_{\text{max}}$, which occurs with probability $1-2^{-t_{\text{max}}}\geq 1-\delta$.
This algorithm to perform this sampling will proceed in two steps. Firstly, it samples a value of $t=1,\dots,t_{\text{max}}$ from the distribution $\frac{2^{-t}}{1-2^{-t_{\text{max}}}}$, which may be done exactly and efficiently. Then, given the sampled $t$, it must sample $I$ with probability $1-c_{t}(2N\mathcal{R})^{t}$ and each $I+(N\mathcal{R})^{-t}\Gamma_{j,t}$ with probability $c_{j,t}(2N\mathcal{R})^{t}$. However, Corollary 3.3 is able to sample each $\Gamma_{j,t}$ with probabilities $p_{j,t}\geq c_{j,t}(2N\mathcal{R})^{t}$, in time. After sampling an operator $\Gamma_{j,t}$, the algorithm can further sample from $\{0,\Gamma_{j,t}\}$ with probabilities $1-\frac{p_{j,t}}{(2N\mathcal{R})^{-t}}$ and $\frac{p_{j,t}}{c_{j,t}(2N\mathcal{R})^{-t}}$. We are now sampling each $\Gamma_{j,t}$ with probability $c_{j,t}(2N\mathcal{R})^{-t}$ and 00 with the remaining probability. Returning $I+(2N\mathcal{R})^{-t}\Gamma_{j,t}$ or $I$ correspondingly gives the desired algorithm.
The above algorithm chooses a value for $t\leq t_{\text{max}}$, and then applies Corollary 3.3 for $t$, which is efficient in $\log\delta^{-1}$ and $n$. Furthermore, Corollary 3.3 guarantees that the number of possible terms $(c_{j,t},\Gamma_{j,t})$ for a given $t$ is at most $(t+1)!\cdot(d+1)^{t}.$ Since we truncate the expansion at $t_{\text{max}}=\lceil\log_{2}(\delta^{-1})\rceil$, the total number of possible sampling outcomes is at most
$\sum_{t=0}^{t_{\text{max}}}(t+1)!\cdot(d+1)^{t}\leq(t_{\text{max}}+1)\cdot(t_{\text{max}}+1)!\cdot(d+1)^{t_{\text{max}}}.$ (55)
Because $\log(\delta^{-1})!$ is quasipolynomial in $\delta^{-1}$, $(t_{\text{max}}+1)!=\text{exp}(\text{polylog}(\delta^{-1}))$. Likewise, $(d+1)^{t_{\text{max}}}=O(\delta^{-\log(d+1)})=\text{poly}(\delta^{-1})$. We conclude that the support of the sampling distribution is $\text{exp}(\text{polylog}(\delta^{-1}))$.
Using the above lemma, we can adapt Algorithm 5.4 to make it efficient to implement, stated below:
Algorithm 5.13 (Pinning one site, efficient)
Input: Subset of terms $S\subseteq[2n]$ for unpinned sites and iteration $j$. Pinned unnormalized Gaussian states $\sigma$ and unpinned state $I+\alpha\Gamma.$ Desired trace distance $\epsilon$ between the sampled state and the Gibbs state, for $\epsilon\geq 0.$
Output: $(j^{\prime},S^{\prime},\sigma^{\prime},\Gamma^{\prime},\alpha^{\prime})$, representing the pinned version of the input states at the next iteration.
  1. 1.
    If $\text{supp}(\Gamma)\cap S$ is nonempty, set $i$ as the minimal index in $\text{supp}(\Gamma)\cap S$. Otherwise, let $i$ be the minimal index in $S$. Define $S^{\prime}=S-\{i\}$. Define
    $M=e^{\beta H^{S^{\prime}/2}}e^{-\beta H^{S/2}}.$
  2. 2.
    Sample $I+\beta_{1}\Lambda_{1}$ for $\beta_{1}=(25\mathcal{R})^{-t_{1}}$ from the probability distribution in Lemma 5.12 with $\delta=\frac{\epsilon}{16n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}$ for $\gamma_{\scriptscriptstyle\mathcal{T}}$ as defined in Lemma 5.10, so that
    $\mathbb{E}\left[I+\beta_{1}\Lambda_{1}\right]=M=e^{\beta H^{S^{\prime}/2}}e^{-\beta H^{S/2}}.$
  3. 3.
    Sample $I+\beta_{2}\Lambda_{2}$ for $\beta_{2}=(25\mathcal{R})^{-t_{2}}$ independently from the same distribution, so that
    $\mathbb{E}\left[I+\beta_{2}\Lambda_{2}\right]=M^{\dagger}.$
  4. 4.
    Set $\sigma^{\prime}\leftarrow\sigma$ and $(\Gamma^{\prime},\alpha^{\prime})$ as:
    1. (a)
      $\bigl{(}\Gamma,(1-\frac{1}{\mathcal{R}})^{-1}\alpha_{j}\bigl{)}$ with probability $1-\frac{1}{\mathcal{R}}$
    2. (b)
      $\left(\Lambda_{1},(6\mathcal{R})\beta_{1}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    3. (c)
      $\left(\Lambda_{2},(6\mathcal{R})\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    4. (d)
      $\left(\Lambda_{1}\Gamma,(6\mathcal{R})\beta_{1}\alpha_{j}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    5. (e)
      $\left(\Gamma\Lambda_{2},(6\mathcal{R})\alpha_{j}\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    6. (f)
      $\left(\Lambda_{1}\Lambda_{2},(6\mathcal{R})\beta_{1}\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
    7. (g)
      $\left(\Lambda_{1}\Gamma\Lambda_{2},(6\mathcal{R})\beta_{1}\alpha_{j}\beta_{2}\right)$ with probability $\frac{1}{6\mathcal{R}}$
  5. 5.
    If $\Gamma^{\prime}$ is not Hermitian:
    • $\Gamma^{\prime}\leftarrow 0,\alpha^{\prime}\leftarrow 0$.
  6. 6.
    If there are two indices $k,l\in S_{j}^{c}$ such that $k,l\in\text{supp}\left(\Gamma^{\prime}\right)$, i.e. $\Gamma^{\prime}=\Gamma^{-}\cdot(i\gamma_{k}\gamma_{l})$:
    1. (a)
      $\Gamma^{\prime}\leftarrow\Gamma^{-}$, $\sigma^{\prime}\leftarrow(I+i\gamma_{k}\gamma_{l})\cdot\sigma^{\prime}$ with probability $\frac{1}{2}$
    2. (b)
      $\Gamma^{\prime}\leftarrow-\Gamma^{-}$, $\sigma^{\prime}\leftarrow(I-i\gamma_{k}\gamma_{l})\cdot\sigma^{\prime}$ with probability $\frac{1}{2}$
  7. 7.
    If $\Gamma^{\prime}\in\{\pm I\}$:
    • $\sigma^{\prime}\leftarrow(I+\alpha^{\prime}\Gamma^{\prime})\sigma^{\prime}$, $\Gamma^{\prime}\leftarrow 0$, and $\alpha^{\prime}\leftarrow 0$.
  8. 8.
    $j\leftarrow j+1$.
Indeed, the rest of the algorithm consists of products of scalars and Majorana strings, which can be encoded efficiently. For this new algorithm, we may now define a new tree $\mathcal{T}^{\prime}$, where again the root node is $(0,S_{0},\sigma_{0},\Gamma_{0},\alpha_{0})$, and the children of a given node are obtained by an application of Algorithm 5.13. Note that $\mathcal{T}^{\prime}$ can be naturally identified as a subtree of $\mathcal{T}$, because the two algorithms are identical except that the sampling distribution in Lemma 5.12 has support contained in that of Lemma 3.2. With this algorithm, we may define a distribution $p^{\prime}$ on the leaves, where $p^{\prime}_{w}$ is the probability of reaching the node $w$ from the root with a repeated application of Algorithm 5.13. Let $\hat{p}^{\prime}_{w}=p^{\prime}_{w}$ for any leaf. We may also define an unnormalized distribution $\hat{q}^{\prime}$ on the leaves by
$\hat{q}^{\prime}_{w}=\text{tr}(\sigma_{w})\hat{p}^{\prime}_{w}.$ (56)
Here, $\hat{q}^{\prime}$ is being defined in direct analogy to $\hat{q}$. Then, $q^{\prime}$ is simply its normalization, i.e.
$q^{\prime}_{v}=\frac{\hat{q}^{\prime}_{v}}{\sum_{w\in\mathcal{L}_{\mathcal{T}^{\prime}}(r)}\hat{q}^{\prime}_{v}}=\frac{\hat{q}^{\prime}_{v}}{\hat{q}^{\prime}_{r}},$ (57)
where $r$ denotes the root node.
Using Lemma 5.12, we can establish a relationship between $p_{v}$ and $p_{v}^{\prime}$. In particular, for any vertex $v$ in the tree $\mathcal{T}^{\prime}$, their difference is at most $Cp_{v}^{\prime}$ for an appropriate constant that is determined by our choice of $\delta$. This bound is obtained by using a similar bound on the conditional distributions of the children of $v$, specifically $\frac{p_{w}^{\prime}}{p_{v}^{\prime}}$ and $\frac{p_{w}}{p_{v}}$.
Lemma 5.14
For the distribution $p^{\prime}$ obtained from Algorithm  5.13, and for any node $v$ and child $w$ in the tree $\mathcal{T}^{\prime}$, $\left|\frac{p_{w}^{\prime}}{p_{v}^{\prime}}-\frac{p_{w}}{p_{v}}\right|\leq\frac{\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\left(\frac{p_{w}^{\prime}}{p_{v}^{\prime}}\right)$. Moreover, for any leaf node $v\in\mathcal{L}_{\mathcal{T}^{\prime}}(r)$, $|p_{v}-p_{v}^{\prime}|\leq\frac{\epsilon}{4(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}p_{v}^{\prime}$.
Proof.
The only difference between Algorithm 5.4 and Algorithm 5.13 is that the sampling in steps 2 and 3 are from different distributions. By Lemma 5.12, the distribution in Algorithm 5.13 is simply the distribution in Algorithm 5.4 conditioned on an event of probability at least $1-\delta$.
There are two such events $A_{1},A_{2}$ that occur in each application of Algorithm 5.13, more specifically in steps 2 and 3. As such, the conditional distribution $\frac{p_{w}^{\prime}}{p_{v}^{\prime}}$ for $w$, i.e. the children of a fixed node $v$, is precisely $\frac{p_{w}}{p_{v}}$ conditioned on the joint event $A_{1}\cap A_{2}$, which happens as an event of probability at least $1-2\delta$ by a union bound. We choose $\delta=\frac{\epsilon}{16n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}.$ It follows that
$\left|\frac{p_{w}}{p_{v}}-\frac{p_{w}^{\prime}}{p_{v}^{\prime}}\right|=\frac{p_{w}^{\prime}}{p_{v}^{\prime}}-\frac{p_{w}}{p_{v}}\leq\frac{p_{w}^{\prime}}{p_{v}^{\prime}}-\left(1-\frac{\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\right)\frac{p_{w}^{\prime}}{p_{v}^{\prime}}=\frac{\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\left(\frac{p_{w}^{\prime}}{p_{v}^{\prime}}\right),$
where the first inequality holds since $\frac{p_{w}^{\prime}}{p_{v}^{\prime}}=\frac{p_{w}/p_{v}}{\mathbb{P}_{p}[A_{1}\cap A_{2}]}\leq\frac{p_{w}/p_{v}}{1-\epsilon/8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}$.
To traverse the entire tree to a leaf, there are $2n$ steps, which means we iteratively apply Algorithm 5.13 $2n$ times. Thus, the distribution $p_{v}^{\prime}$ over the leaves $v$ is $p_{v}$ conditioned on an event of probability at least $1-2n\cdot 2\delta=1-\frac{\epsilon}{4(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}$ by a union bound as well. For any leaf node $v\in\mathcal{L}_{\mathcal{T}^{\prime}}(r)$, it follows by the same reasoning as the above that
$|p_{v}-p_{v}^{\prime}|\leq\frac{\epsilon}{4(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}p_{v}^{\prime}.$
For our new tree $\mathcal{T}^{\prime}$, there is also a natural bound on the number of children of any node.
Lemma 5.15
The number of children of any node in $\mathcal{T}^{\prime}$ is at most $k$ for $\log(k)=\text{poly}(n,\log(\epsilon^{-1}))$.
Proof.
The sampling algorithm in Algorithm 5.4 involves three random choices. The first two are calls to the efficient sampling algorithm from Lemma 5.12, in step 2 and 3 of the algorithm respectively, which has support with size $\text{exp}(\text{poly}(\log(\delta^{-1})))$. We will use a more crude bound of $\text{exp}(\text{poly}(\delta^{-1}))$ for the support to simplify the choice of $\delta$ below. We choose $\delta=\frac{\epsilon}{16n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}$, from which we get the number of possible outcomes in the form of $(t_{i},I+\beta_{t_{i}}\Lambda_{t_{i}})$ for each of these two calls is bounded by $\text{exp}(\text{poly}(n,\log(\epsilon^{-1}))).$ The next two calls are random choices from seven choices of $(\Gamma,\sigma,\alpha)$ in step 4 and two options of $(\Gamma,\sigma,\alpha)$ in step 6, giving an additional factor of 14, which is constant. This gives us the final number of children of at most $k=\text{exp}(\text{poly}(n,\log(\epsilon^{-1}))).$
We will now establish two important results that relate distributions on $\mathcal{T}$ to those on $\mathcal{T}^{\prime}$, presented in Lemma 5.16 and Lemma 5.17 below, which will each be important in the proof of Theorem 5.11. Lemma 5.16 is the most technical, bounding the difference between $\frac{\hat{q}_{v}}{\hat{p}_{v}}$ and $\frac{\hat{q}^{\prime}_{v}}{\hat{p}^{\prime}_{v}}$. We have previously seen $\frac{\hat{q}_{v}}{\hat{p}_{v}}$ to be the expected value of $\text{tr}(\sigma_{w})$ for leaves $w$ below $v$ under the distribution $p_{w}$, so intuitively the lemma establishes how well the primed distribution approximates these conditional expectations. It has two roles: it will allow us to relate $\hat{q}_{r}$ to $\hat{q}^{\prime}_{r}$, and bound the multiplicative distance between $\frac{\hat{q}_{v}^{\prime}}{\hat{p}_{v}^{\prime}}$ and $\frac{\hat{q}_{w}^{\prime}}{\hat{p}_{w}^{\prime}}$ for a node $v$ and its child $w$ (to satisfy condition (1)) .
Lemma 5.16
For any node $v\in\mathcal{T}^{\prime}$,
$\left|\frac{\hat{q}^{\prime}_{v}}{\hat{p}^{\prime}_{v}}-\frac{\hat{q}_{v}}{\hat{p}_{v}}\right|\leq\frac{(2n-j_{v})(\gamma_{\scriptscriptstyle\mathcal{T}}+1)\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\max\left(\frac{\hat{q}_{v}^{\prime}}{\hat{p}_{v}^{\prime}},\frac{\hat{q}_{v}}{\hat{p}_{v}}\right).$
In particular,
$\left|\log\left(\frac{\hat{q}^{\prime}_{v}}{\hat{p}^{\prime}_{v}}\right)-\log\left(\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)\right|\leq\log\left(1+\frac{(2n-j_{v})(\gamma_{\scriptscriptstyle\mathcal{T}}+1)\epsilon}{4n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\right).$
Proof.
We will prove this reduction by descending strong induction on the depth of $v$, with the base case of $j_{v}=2n$. In this case, by definition $\frac{\hat{q}_{v}^{\prime}}{\hat{p}_{v}^{\prime}}=\frac{\hat{q}_{v}}{\hat{p}_{v}}=\text{tr}(\sigma_{v})$, so the result is immediate. Now see that we may write $\frac{\hat{q}^{\prime}_{v}}{\hat{p}^{\prime}_{v}}$ as
$\frac{\hat{q}^{\prime}_{v}}{\hat{p}^{\prime}_{v}}=\sum_{w\in\mathcal{C}_{\mathcal{T}^{\prime}}(v)}\left(\frac{\hat{q}^{\prime}_{w}}{\hat{p}^{\prime}_{w}}\right)\cdot\left(\frac{\hat{p}^{\prime}_{w}}{\hat{p}^{\prime}_{v}}\right)=\sum_{w\in\mathcal{C}_{\mathcal{T}^{\prime}}(v)}\left(\frac{\hat{q}^{\prime}_{w}}{\hat{p}^{\prime}_{w}}\right)\cdot\left(\frac{\hat{p}_{w}}{\hat{p}_{v}}\right)+\sum_{w\in\mathcal{C}_{\mathcal{T}^{\prime}}(v)}\left(\frac{\hat{q}^{\prime}_{w}}{\hat{p}^{\prime}_{w}}\right)\left(\frac{\hat{p}^{\prime}_{w}}{\hat{p}^{\prime}_{v}}-\frac{\hat{p}_{w}}{\hat{p}_{v}}\right),$
Likewise,
$\frac{\hat{q}_{v}}{\hat{p}_{v}}=\sum_{w\in\mathcal{C}_{\mathcal{T}}(v)}\left(\frac{\hat{q}_{w}}{\hat{p}_{w}}\right)\cdot\left(\frac{\hat{p}_{w}}{\hat{p}_{v}}\right).$
The proof of Lemma 5.14 states that the distribution $\frac{p^{\prime}_{w}}{p^{\prime}_{v}}$ is the distribution $\frac{p_{w}}{p_{v}}$ conditioned on an event with probability at least $1-\frac{\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}$.
By induction, we assume that for all children nodes $w\in\mathcal{C}_{\mathcal{T}^{\prime}}(v)$ (which necessarily satisfy $j_{w}=j_{v}+1$), the following holds:
$\left|\frac{\hat{q}_{w}^{\prime}}{\hat{p}_{w}^{\prime}}-\frac{\hat{q}_{w}}{\hat{p}_{w}}\right|\leq\left(\frac{(2n-j_{w})(\gamma_{\scriptscriptstyle\mathcal{T}}+1)\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\right)\max\left(\frac{\hat{q}_{v}^{\prime}}{\hat{p}_{v}^{\prime}},\frac{\hat{q}_{v}}{\hat{p}_{v}}\right).$
The base case for leaf nodes is clear because they are precisely equal, both $\text{tr}(\sigma_{w})$. Now see that
$\left|\frac{\hat{q}_{v}^{\prime}}{\hat{p}_{v}^{\prime}}-\frac{\hat{q}_{v}}{\hat{p}_{v}}\right|$
$=\left|\sum_{w\in\mathcal{C}_{\mathcal{T}^{\prime}}(v)}\left(\frac{\hat{q}_{w}^{\prime}}{\hat{p}_{w}^{\prime}}-\frac{\hat{q}_{w}}{\hat{p}_{w}}\right)\cdot\left(\frac{\hat{p}_{w}}{\hat{p}_{v}}\right)+\sum_{w\in\mathcal{C}_{\mathcal{T}^{\prime}}(v)}\left(\frac{\hat{q}^{\prime}_{w}}{\hat{p}^{\prime}_{w}}\right)\left(\frac{\hat{p}^{\prime}_{w}}{\hat{p}^{\prime}_{v}}-\frac{\hat{p}_{w}}{\hat{p}_{v}}\right)-\sum_{w\in\mathcal{C}_{\mathcal{T}}(v)\backslash\mathcal{C}_{\mathcal{T}^{\prime}}(v)}\left(\frac{\hat{q}_{w}}{\hat{p}_{w}}\right)\cdot\left(\frac{\hat{p}_{w}}{\hat{p}_{v}}\right)\right|$
$\leq\sum_{w\in\mathcal{C}_{\mathcal{T}}^{\prime}(v)}\left|\frac{\hat{q}_{w}^{\prime}}{\hat{p}_{w}^{\prime}}-\frac{\hat{q}_{w}}{\hat{p}_{w}}\right|\cdot\left(\frac{\hat{p}_{w}}{\hat{p}_{v}}\right)+\frac{\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\left(\frac{\hat{q}_{v}^{\prime}}{\hat{p}_{v}^{\prime}}\right)+\frac{\gamma_{\scriptscriptstyle\mathcal{T}}\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\left(\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)$
$\leq\frac{(2n-j_{v}-1)(\gamma_{\scriptscriptstyle\mathcal{T}}+1)\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\max\left(\frac{\hat{q}_{v}^{\prime}}{\hat{p}_{v}^{\prime}},\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)+\frac{\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\left(\frac{\hat{q}_{v}^{\prime}}{\hat{p}_{v}^{\prime}}\right)+\frac{\gamma_{\scriptscriptstyle\mathcal{T}}\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\left(\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)$
$\leq\frac{(2n-j_{v})(\gamma_{\scriptscriptstyle\mathcal{T}}+1)\epsilon}{8n(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\max\left(\frac{\hat{q}_{v}^{\prime}}{\hat{p}_{v}^{\prime}},\frac{\hat{q}_{v}}{\hat{p}_{v}}\right),$
where we use Lemma 5.14 in the first inequality, as well as the fact that $\frac{\hat{q}_{w}}{\hat{p}_{w}}\leq\gamma_{\scriptscriptstyle\mathcal{T}}\frac{\hat{q}_{v}}{\hat{p}_{v}}$ for any child $w$ of $v$. To obtain the second statement, note that since $\frac{\epsilon(\gamma_{\scriptscriptstyle\mathcal{T}}+1)}{4(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\leq\frac{\epsilon}{4}\leq\frac{1}{2}$, explicit calculation shows us that the first statement implies that
$\left|\log\left(\frac{\hat{q}^{\prime}_{v}}{\hat{p}^{\prime}_{v}}\right)-\log\left(\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)\right|\leq\log\left(1+2\left(\frac{\epsilon(\gamma_{\scriptscriptstyle\mathcal{T}}+1)}{4(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\right)\right)=\log\left(1+\frac{\epsilon(\gamma_{\scriptscriptstyle\mathcal{T}}+1)}{2(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\right).$
We are now ready to prove that sampling from $q^{\prime}_{v}$ amount to sampling from $q_{v}$, up to a total variation distance of $\frac{\epsilon}{2}$. The proof of this lemma reduces to the previous bound on $|p_{v}-p_{v}^{\prime}|$, and the technical result of Lemma 5.16 that can bound the difference of $\hat{q}_{r}^{\prime}$ and $\hat{q}_{r}$. The constant $\delta$ was chosen precisely to obtain the following bound.
Lemma 5.17
The total variation distance of the distributions $q_{v}$ and $q^{\prime}_{v}$ is bounded by $\frac{\epsilon}{2}$.
Proof.
The total variation distance of the two distributions may be written as
$\text{TV}(q,q^{\prime})$ $=\frac{1}{2}\sum_{v\in\mathcal{L}_{\mathcal{T}(r)}}|q_{v}-q_{v}^{\prime}|$
$=\frac{1}{2}\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}|q_{v}-q_{v}^{\prime}|+\frac{1}{2}\sum_{v\in\mathcal{L}_{\mathcal{T}(r)}\backslash\mathcal{L}_{\mathcal{T^{\prime}}(r)}}|q_{v}-q_{v}^{\prime}|$
$=\frac{1}{2}\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}|q_{v}-q_{v}^{\prime}|+\frac{1}{2}\sum_{v\in\mathcal{L}_{\mathcal{T}(r)}\backslash\mathcal{L}_{\mathcal{T^{\prime}}(r)}}q_{v}$
$\leq\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}|q_{v}-q_{v}^{\prime}|,$
where the second equality holds since $\hat{q}^{\prime}_{v}=0$ for nodes $v\notin\mathcal{L}_{\mathcal{T}^{\prime}}(r)$, and the inequality holds because
$\sum_{v\in\mathcal{L}_{\mathcal{T}(r)}\backslash\mathcal{L}_{\mathcal{T^{\prime}}(r)}}q_{v}=1-\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}}(r)}q_{v}=\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}}(r)}q^{\prime}_{v}-q_{v}\leq\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}}(r)}|q_{v}-q^{\prime}_{v}|.$
Now, it follows that, denoting the root of $\mathcal{T}^{\prime}$ by $r$,
$\text{TV}(q,q^{\prime})$ $\leq\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}|q_{v}-q_{v}^{\prime}|$
$=\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}\left|\frac{\hat{q}_{v}}{\hat{q}_{r}}-\frac{\hat{q}_{v}^{\prime}}{\hat{q}_{r}^{\prime}}\right|$
$\leq\frac{1}{\hat{q}_{r}^{\prime}}\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}\left|\hat{q}_{v}-\hat{q}_{v}^{\prime}\right|+\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}\left|\frac{\hat{q}_{v}}{\hat{q}_{r}}-\frac{\hat{q}_{v}}{\hat{q}_{r}^{\prime}}\right|$
$=\frac{1}{\hat{q}_{r}^{\prime}}\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}\text{tr}(\sigma_{v})|p_{v}-p_{v}^{\prime}|+\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}\frac{\hat{q}_{v}}{\hat{q}_{r}\hat{q}_{r}^{\prime}}\left|\hat{q}_{r}^{\prime}-\hat{q}_{r}\right|$
$\leq\frac{1}{\hat{q}_{r}^{\prime}}\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}}(r)}\text{tr}(\sigma_{v})|p_{v}-p_{v}^{\prime}|+\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}\frac{\hat{q}_{v}}{\hat{q}_{r}}\left(\frac{\epsilon(\gamma_{\scriptscriptstyle\mathcal{T}}+1)}{2(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\right)$
$\leq\frac{\epsilon}{4(\gamma_{\scriptscriptstyle\mathcal{T}}+2)\hat{q}_{r}^{\prime}}\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}}(r)}p^{\prime}_{v}\text{tr}(\sigma_{v})+\frac{\epsilon(\gamma_{\scriptscriptstyle\mathcal{T}}+1)}{2(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}$
$=\frac{\epsilon}{4(\gamma_{\scriptscriptstyle\mathcal{T}}+2)\hat{q}_{r}^{\prime}}\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}}(r)}\hat{q}^{\prime}_{v}+\frac{\epsilon(\gamma_{\scriptscriptstyle\mathcal{T}}+1)}{2(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}$
$<\frac{\epsilon}{2}$
as desired. Here, the second inequality comes from applying a direct triangle inequality, and the third inequality comes from Lemma 5.16, applied to the root node $r$. In particular, since $\hat{p}_{r}=\hat{p}_{r}^{\prime}=1$, we have that $\left|\log(\hat{q}^{\prime}_{r})-\log(\hat{q}_{r})\right|\leq\log(1+\frac{\epsilon(\gamma_{\scriptscriptstyle\mathcal{T}}+1)}{2(\gamma_{\scriptscriptstyle\mathcal{T}}+2)})$, which implies that $|\hat{q}_{r}-\hat{q}^{\prime}_{r}|\leq\frac{\epsilon(\gamma_{\scriptscriptstyle\mathcal{T}}+1)}{2(\gamma_{\scriptscriptstyle\mathcal{T}}+2)}\hat{q}^{\prime}_{r}$. The last inequality comes from applying Lemma 5.14 and the fact that $\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}(r)}}\frac{\hat{q}_{v}}{\hat{q}_{r}}\leq 1.$ We further note that $\sum_{v\in\mathcal{L}_{\mathcal{T}^{\prime}}(r)}\hat{q}^{\prime}_{v}=\hat{q}^{\prime}_{r}$ to arrive at the final expression. ∎
Having developed approximate analogs of the original distributions on $\mathcal{T}$, we can prove Theorem 5.11.
Theorem 5.11.
This proof will proceed in two steps. Firstly, we will show that all four conditions of Theorem 5.5 are satisfied for the distributions $p^{\prime}$ and $q^{\prime}$. Then, we will show that the ability to sample from $q^{\prime}$ yields the desired Gibbs sampling algorithm.
We will establish each condition of $p^{\prime}$ and $q^{\prime}$ necessary for Theorem 5.5. We also have that condition (2) is easily satisfied. Indeed, by definition, for any leaf node $v\in\mathcal{T}^{\prime}$
$\frac{\hat{q}_{v}^{\prime}}{\hat{p}^{\prime}_{v}}=\text{tr}(\sigma_{v}),$
which can be exactly calculated given $v$. Condition (3) is also satisfied. Indeed, for any internal node $v$ and a child $v_{k}$, $\frac{p^{\prime}_{v_{k}}}{p^{\prime}_{v}}$ is the conditional probability of reaching $v_{k}$ by repeated application of Algorithm 5.13 given that the algorithm reaches $v$. A single application of Algorithm 22 therefore suffices to sample the children of $v$ according to this probability distribution. Now notice that both conditions (1) and (4) follow from the statement of Lemma 5.16, which implies that
$\left|\log\left(\frac{\hat{q}^{\prime}_{v}}{\hat{p}^{\prime}_{v}}\right)-\log\left(\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)\right|\leq\log\left(1+\frac{\epsilon}{2}\right).$
As we have already demonstrated an efficient algorithm to calculate $\frac{\hat{q}_{v}}{\hat{p}_{v}}$ up to constant multiplicative error, this statement shows that the same algorithm calculates $\frac{\hat{q}_{v}^{\prime}}{\hat{p}_{v}^{\prime}}$ up to constant multiplicative error. Condition (1) is therefore satisfied. Meanwhile, condition (4) can be obtained from the triangle inequality, since for a node $v$ and a child $w$,
$\left|\log\left(\frac{\hat{q}^{\prime}_{v}}{\hat{p}^{\prime}_{v}}\right)-\log\left(\frac{\hat{q}^{\prime}_{w}}{\hat{p}^{\prime}_{w}}\right)\right|$ $\leq\left|\log\left(\frac{\hat{q}^{\prime}_{v}}{\hat{p}^{\prime}_{v}}\right)-\log\left(\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)\right|+\left|\log\left(\frac{\hat{q}_{v}}{\hat{p}_{v}}\right)-\log\left(\frac{\hat{q}_{w}}{\hat{p}_{w}}\right)\right|$
$+\left|\log\left(\frac{\hat{q}^{\prime}_{w}}{\hat{p}^{\prime}_{w}}\right)-\log\left(\frac{\hat{q}_{w}}{\hat{p}_{w}}\right)\right|$
$\leq 2\log\left(1+\frac{\epsilon}{2}\right)+\log(\gamma_{\scriptscriptstyle\mathcal{T}}),$
where the middle term is at most $\log(\gamma_{\scriptscriptstyle\mathcal{T}})$ since condition (4) is satisfied for $\hat{p}$ and $\hat{q}$. We may therefore apply Theorem 5.5 and obtain a $\text{poly}(n,\log(\epsilon^{-1}),\log(k))$ algorithm to sample a distribution $q^{\prime\prime}$ that is $q^{\prime}$ up to $\frac{\epsilon}{2}$ total variation distance. The number of children of each node was bounded in Lemma 5.15 so that $\log(k)$ is polynomial in $n$. By Lemma 5.17, the total variation distance between $q^{\prime\prime}$ and $q$ is at most $\epsilon$. It follows that
$\left\lVert\sum_{v\in\mathcal{L}_{\mathcal{T}}(r)}q^{\prime\prime}_{v}\frac{\sigma_{v}}{\text{tr}(\sigma_{v})}-\rho_{\beta}\right\rVert_{\text{tr}}$ $=\frac{1}{2}\left\lVert\sum_{v\in\mathcal{L}_{\mathcal{T}}(r)}q^{\prime\prime}_{v}\frac{\sigma_{v}}{\text{tr}(\sigma_{v})}-\rho_{\beta}\right\rVert_{1}$
$=\frac{1}{2}\left\lVert\sum_{v\in\mathcal{L}_{\mathcal{T}}(r)}q^{\prime\prime}_{v}\frac{\sigma_{v}}{\text{tr}(\sigma_{v})}-\sum_{v\in\mathcal{L}_{\mathcal{T}}(r)}q_{v}\frac{\sigma_{v}}{\text{tr}(\sigma_{v})}\right\rVert_{1}$
$\leq\frac{1}{2}\sum_{v\in\mathcal{L}_{\mathcal{T}}(r)}|q^{\prime\prime}_{v}-q_{v}|$
$\leq\epsilon,$
giving the desired trace distance bound.
6 Conclusion
Motivated by recent progress on the lack of entanglement in high-temperature Gibbs states of quantum spin systems [5], in this work we showed that high-temperature Gibbs states of fermionic systems are probabilistic mixtures of fermionic Gaussian states. Fermionic Gaussian states are completely determined by their correlation matrices, which enables efficient classical representation of these states. Based on this structural result, we then proposed an efficient algorithm to sample from this distribution of fermionic Gaussian states, thus enabling efficient computation of observable expectation values and sampling from the Gibbs states.
Besides techniques developed in [5], our results rely on a new adaptive pairing procedure for Majorana modes. While for quantum spin systems locality of the operators can imply commutativity, such connection is only possible in the fermionic case with an appropriate pairing.
Our result sharpens the the boundary between quantum advantage and classical simulability for fermionic systems. Recent results on the mixing time of weakly interacting fermionic systems shows that at any constant temperature, the corresponding Gibbs states can be efficiently prepared on a quantum computer [20, 21]. Our result shows that in the high-temperature regime, this mixing time result does not lead to a super-polynomial quantum advantage. However, for a temperature below the threshold considered in this work, the existence of significant quantum advantage remains an open problem. Given the importance of the structural result in the design of the algorithm, it is natural to ask whether weakly interacting fermionic Gibbs states can also be expressed as mixtures of fermionic Gaussian states.
The connection between fermionic Gaussian mixtures and simulability was also explored in the simulation of fermionic circuits [26]. Given the prominent role played by this structure in the simulability of fermionic systems, it is of interest to explore ways to test whether a fermionic quantum state is a mixture of Gaussian states or not. This question can be asked in at least two different settings: Firstly, given a quantum state that is the ground state or the Gibbs state of a non-quadratic fermionic Hamiltonian, can we distinguish whether it is a mixture of fermionic Gaussian states? Secondly, if we have an unknown quantum state on an experimental fermionic system, what are the sample and computational complexities of answering such a question? Answering these questions can help us better understand the classical simulability of fermionic systems and the boundary of quantum advantage.
Acknowledgments
We thank Ewin Tang, Ainesh Bakshi, and Robbie King for helpful discussions. Y.C. is supported by the Mellon Mays Undergraduate Fellowship. J.J. is supported by MURI Grant FA9550-18-1-0161, NSF CAREER award CCF-2048204, and the IQIM, an NSF Physics Frontiers Center (NSF Grant PHY-1125565).